RTEMS Shell User’s Guide

Edition 4.10.0, for RTEMS 4.10.0

28 February 2008

On-Line Applications Research Corporation

On-Line Applications Research Corporation TEXinfo 2009-08-14.15

COPYRIGHT \begin{matrix}c \\ ◯ \end{matrix} 1988 - 2008. On-Line Applications Research Corporation (OAR).

The authors have used their best efforts in preparing this material. These efforts include the development, research, and testing of the theories and programs to determine their effectiveness. No warranty of any kind, expressed or implied, with regard to the software or the material contained in this document is provided. No liability arising out of the application or use of any product described in this document is assumed. The authors reserve the right to revise this material and to make changes from time to time in the content hereof without obligation to notify anyone of such revision or changes.

The RTEMS Project is hosted at http://www.rtems.com. Any inquiries concerning RTEMS, its related support components, its documentation, or any custom services for RTEMS should be directed to the contacts listed on that site. A current list of RTEMS

Support Providers is at http://www.rtems.com/support.html.

Table of Contents

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

1 Configuration and Initialization . . . . . . . . . . . . . . . 3

1.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.2.1 Customizing the Command Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2.2 Adding Custom Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.3 Initialization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.3.1 Attached to a Serial Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.3.2 Attached to a Socket . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

1.4 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.4.1 rtems shell init - initialize the shell . . . . . . . . . . . . . . . . . . . . . . . . 6

2 General Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2 Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.2.1 alias - add alias for an existing command . . . . . . . . . . . . . . . . . . 8 2.2.2 date - print or set current date and time . . . . . . . . . . . . . . . . . . . 9 2.2.3 echo - produce message in a shell script . . . . . . . . . . . . . . . . . . . 10 2.2.4 sleep - delay for a specified amount of time . . . . . . . . . . . . . . . 12 2.2.5 id - show uid gid euid and egid . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.2.6 tty - show ttyname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.2.7 whoami - print effective user id . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.2.8 getenv - print environment variable . . . . . . . . . . . . . . . . . . . . . . . 16 2.2.9 setenv - set environment variable . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.2.10 unsetenv - unset environment variable . . . . . . . . . . . . . . . . . . . 18 2.2.11 time - time command execution . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.2.12 logoff - logoff from the system . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.2.13 rtc - RTC driver configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.2.14 exit - exit the shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

3 File and Directory Commands . . . . . . . . . . . . . . . 23

3.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.2 Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

3.2.1 umask - set file mode creation mask . . . . . . . . . . . . . . . . . . . . . . 24 3.2.2 cp - copy files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.2.3 mv - move files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.2.4 pwd - print work directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.2.5 ls - list files in the directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.2.6 chdir - change the current directory. . . . . . . . . . . . . . . . . . . . . . . 33 3.2.7 mkdir - create a directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 3.2.8 rmdir - remove empty directories . . . . . . . . . . . . . . . . . . . . . . . . . 36 3.2.9 ln - make links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.2.10 mknod - make device special file. . . . . . . . . . . . . . . . . . . . . . . . . 39 3.2.11 chroot - change the root directory . . . . . . . . . . . . . . . . . . . . . . . 41 3.2.12 chmod - change permissions of a file . . . . . . . . . . . . . . . . . . . . . 42 3.2.13 cat - display file contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 3.2.14 rm - remove files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 3.2.15 mount - mount disk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 3.2.16 unmount - unmount disk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 3.2.17 blksync - sync the block driver . . . . . . . . . . . . . . . . . . . . . . . . . . 49 3.2.18 dd - convert and copy a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 3.2.19 hexdump - ascii/dec/hex/octal dump . . . . . . . . . . . . . . . . . . . . 55 3.2.20 fdisk - format disk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 3.2.21 dir - alias for ls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 3.2.22 mkrfs - format RFS file system . . . . . . . . . . . . . . . . . . . . . . . . . . 61 3.2.23 debugrfs - debug RFS file system . . . . . . . . . . . . . . . . . . . . . . . . 63 3.2.24 cd - alias for chdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

4 Memory Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . 67

4.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 4.2 Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

4.2.1 mdump - display contents of memory . . . . . . . . . . . . . . . . . . . . . 68 4.2.2 wdump - display contents of memory (word) . . . . . . . . . . . . . . 69 4.2.3 medit - modify contents of memory . . . . . . . . . . . . . . . . . . . . . . . 70 4.2.4 mfill - file memory with pattern . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.2.5 mmove - move contents of memory . . . . . . . . . . . . . . . . . . . . . . . 73 4.2.6 malloc - obtain information on C program heap. . . . . . . . . . . 74

5 RTEMS Specific Commands. . . . . . . . . . . . . . . . . . 77

5.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 5.2 Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

5.2.1 halt - Shutdown the system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 5.2.2 cpuuse - print or reset per thread cpu usage . . . . . . . . . . . . . . 79 5.2.3 stackuse - print per thread stack usage. . . . . . . . . . . . . . . . . . . . 81 5.2.4 perioduse - print or reset per period usage . . . . . . . . . . . . . . . . 82 5.2.5 wkspace - display information on executive workspace. . . . . 84 5.2.6 config - show the system configuration. . . . . . . . . . . . . . . . . . . . 86 5.2.7 itask - list init tasks for the system . . . . . . . . . . . . . . . . . . . . . . . 87 5.2.8 extension - display information about extensions . . . . . . . . . . 88 5.2.9 task - display information about tasks . . . . . . . . . . . . . . . . . . . . 89 5.2.10 queue - display information about message queues . . . . . . . 90 5.2.11 sema - display information about semaphores . . . . . . . . . . . . 91 5.2.12 region - display information about regions . . . . . . . . . . . . . . . 92 5.2.13 part - display information about partitions . . . . . . . . . . . . . . 93 5.2.14 object - display information about rtems objects . . . . . . . . . 94 5.2.15 driver - display the rtems device driver table. . . . . . . . . . . . . 95 5.2.16 dname - displays information about named drivers . . . . . . . 96 5.2.17 pthread - display information about POSIX threads . . . . . 97

6 Network Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . 99

6.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 6.2 Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

6.2.1 netstats - obtain network statistics . . . . . . . . . . . . . . . . . . . . . . 100 6.2.2 ifconfig - configure a network interface . . . . . . . . . . . . . . . . . . . 103 6.2.3 route - show or manipulate the ip routing table . . . . . . . . . . 105

Function and Variable Index . . . . . . . . . . . . . . . . . . . . 107

Concept Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

Command Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

Preface

Real-time embedded systems vary widely based upon their operational and maintenance requirements. Some of these systems provide ways for the user or developer to interact with them. This interaction could be used for operational, diagnostic, or configuration purposes. The capabilities described in this manual are those provided with RTEMS to provide a command line interface for user access. Some of these commands will be familiar as standard POSIX utilities while others are RTEMS specific or helpful in debugging and analyzing an embedded system. As a simple example of the powerful and very familiar capabilities that the RTEMS Shell provides to an application, consider the following example which hints at some of the capabilities available:

      Welcome to rtems-4.10.0(SPARC/w/FPU/sis)
      COPYRIGHT (c) 1989-2011.
      On-Line Applications Research Corporation (OAR).
      Login into RTEMS
      login: rtems
      Password:
      RTEMS SHELL (Ver.1.0-FRC):/dev/console. Feb 28 2008. 'help' to list commands.
      SHLL [/] $ cat /etc/passwd
      root:*:0:0:root::/:/bin/sh
      rtems:*:1:1:RTEMS Application::/:/bin/sh
      tty:!:2:2:tty owner::/:/bin/false
      SHLL [/] $ ls /dev
      -rwxr-xr-x   1  rtems   root           0 Jan 01 00:00 console
      -rwxr-xr-x   1   root   root           0 Jan 01 00:00 console_b
      2 files 0 bytes occupied
      SHLL [/] $ stackuse
      Stack usage by thread
          ID      NAME    LOW          HIGH     CURRENT     AVAILABLE     USED
      0x09010001  IDLE 0x023d89a0 - 0x023d99af 0x023d9760      4096        608
      0x0a010001  UI1  0x023d9f30 - 0x023daf3f 0x023dad18      4096       1804
      0x0a010002  SHLL 0x023db4c0 - 0x023df4cf 0x023de9d0     16384       6204
      0xffffffff  INTR 0x023d2760 - 0x023d375f 0x00000000      4080        316
      SHLL [/] $ mount -L
      File systems: msdos
      SHLL [/] $

In the above example, the user rtems logs into a SPARC based RTEMS system. The first command is cat /etc/passwd. This simple command lets us know that this application is running the In Memory File System (IMFS) and that the infrastructure has provided dummy entries for /etc/passwd and a few other files. The contents of /etc/passwd let us know that the user could have logged in as root. In fact, the root user has more permissions than rtems who is not allowed to write into the filesystem.

The second command is ls /dev which lets us know that RTEMS has POSIX-style device nodes which can be accesses through standard I/O function calls.

The third command executed is the RTEMS specific stackuse which gives a report on the stack usage of each thread in the system. Since stack overflows are a common error in deeply embedded systems, this is a surprising simple, yet powerful debugging aid.

Finally, the last command, mount -L hints that RTEMS supports a variety of mountable filesystems. With support for MS-DOS FAT on IDE/ATA and Flash devices as well as network-based filesystens such as NFS and TFTP, the standard free RTEMS provides a robuse infrastructure for embedded applications.

This manual describes the RTEMS Shell and its command set. In our terminology, the Shell is just a loop reading user input and turning that input into commands with argument. The Shell provided with RTEMS is a simple command reading loop with limited scripting capabilities. It can be connected to via a standard serial port or connected to the RTEMS telnetd server for use across a network.

Each command in the command set is implemented as a single subroutine which has a main-style prototype. The commands interpret their arguments and operate upon stdin, stdout, and stderr by default. This allows each command to be invoked independent of the shell.

The described separation of shell from commands from communications mechanism was an important design goal. At one level, the RTEMS Shell is a complete shell environment providing access to multiple POSIX compliant filesystems and TCP/IP stack. The subset of capabilities available is easy to configure and the standard Shell can be logged into from either a serial port or via telnet. But at another level, the Shell is a large set of components which can be integrated into the user’s developed command interpreter. In either case, it is trivial to add custom commands to the command set available.

Acknowledgements

The Institute of Electrical and Electronics Engineers, Inc and The Open Group, have given us permission to reprint portions of their documentation.

Portions of this text are reprinted and reproduced in electronic form from IEEE Std 1003.1, 2004 Edition, Standard for Information Technology Operating System Interface (POSIX), The Open Group Base Specifications Issue 6, Copyright 2001-2004 by the Institute of Electrical and Electronics Engineers, Inc and The Open Group. In the event of any discrepancy between this version and the original IEEE and The Open Group Standard, the original IEEE and The Open Group Standard is the referee document. The original Standard can be obtained online at http://www.opengroup.org/unix/online.html

     (http://www.opengroup.org/unix/online.html).

This notice shall appear on any product containing this material.

1 Configuration and Initialization

1.1 Introduction

This chapter provides information on how the application configures and intializes the RTEMS shell.

1.2 Configuration

The command set available to the application is user configurable. It is configured using a mechanism similar to the confdefs.h mechanism used to specify application configuration.

In the simplest case, if the user wishes to configure a command set with all commands available that are neither filesystem management (e.g. mounting, formating, etc.) or network related, then the following is all that is required:

      #define CONFIGURE_SHELL_COMMANDS_INIT
      #define CONFIGURE_SHELL_COMMANDS_ALL
      #include <rtems/shellconfig.h>

In a slightly more complex example, if the user wishes to include all networking commands as well as support for mounting MS-DOS and NFS filesystems, then the following is all that is required:

      #define CONFIGURE_SHELL_COMMANDS_INIT
      #define CONFIGURE_SHELL_COMMANDS_ALL
      #define CONFIGURE_SHELL_MOUNT_MSDOS
      #define CONFIGURE_SHELL_MOUNT_NFS
      #include <rtems/shellconfig.h>

1.2.1 Customizing the Command Set

The user can configure specific command sets by either building up the set from individual commands or starting with a complete set and disabling individual commands. Each command has two configuration macros associated with it.

CONFIGURE_SHELL_COMMAND_XXX

Each command has a constant of this form which is defined when building a command set by individually enabling specific commands.

CONFIGURE_SHELL_NO_COMMAND_XXX

In contrast, each command has a similar command which is defined when the application is configuring a command set by disabling specific commands in the set.

1.2.2 Adding Custom Commands

One of the design goals of the RTEMS Shell was to make it easy for a user to add custom commands specific to their application. We believe this design goal was accomplished. In order to add a custom command, the user is required to do the following:

• Provide a main-style function which implements the command. If that command function uses a getopt related function to parse arguments, it MUST use the reentrant form.

• Provide a command definition structure of type rtems_shell_cmd_t.

• Configure that command using the CONFIGURE_SHELL_USER_COMMANDS macro.

Custom aliases are configured similarly but the user only provides an alias definition structure of type rtems_shell_alias_t and configures the alias via the CONFIGURE_SHELL_

USER_ALIASES macro.

In the following example, we have implemented a custom command named usercmd which simply prints the arguments it was passed. We have also provided an alias for usercmd

named userecho.
      #include <rtems/shell.h>
      int main_usercmd(int argc, char **argv)
      {
        int i;
        printf( "UserCommand: argc=%d\n", argc );
        for (i=0 ; i<argc ; i++ )
          printf( "argv[%d]= %s\n", i, argv[i] );
        return 0;
      }
      rtems_shell_cmd_t Shell_USERCMD_Command = {
        "usercmd",                  /* name */
        "usercmd n1 [n2 [n3...]]",  /* usage */
        "user",                     /* topic */
        main_usercmd,               /* command */
        NULL,                       /* alias */
        NULL                        /* next */
      };
      rtems_shell_alias_t Shell_USERECHO_Alias = {
        "usercmd",                  /* command */
        "userecho"                  /* alias */
      };
      #define CONFIGURE_SHELL_USER_COMMANDS &Shell_USERCMD_Command
      #define CONFIGURE_SHELL_USER_ALIASES &Shell_USERECHO_Alias
      #define CONFIGURE_SHELL_COMMANDS_INIT
      #define CONFIGURE_SHELL_COMMANDS_ALL
      #define CONFIGURE_SHELL_MOUNT_MSDOS
      #include <rtems/shellconfig.h>

Notice in the above example, that the user wrote the main for their command (e.g. main_ usercmd) which looks much like any other main(). They then defined a rtems_shell_cmd_t structure named Shell_USERCMD_Command which describes that command. This command definition structure is registered into the static command set by defining CONFIGURE_SHELL_

USER_COMMANDS to &Shell_USERCMD_Command.

Similarly, to add the userecho alias, the user provides the alias definition structure

named Shell_USERECHO_Alias and defines CONFIGURE_SHELL_USER_ALIASES to configure

the alias.

The user can configure any number of commands and aliases in this manner.

1.3 Initialization

The shell may be easily attached to a serial port or to the telnetd server. This section describes how that is accomplished.

1.3.1 Attached to a Serial Port

Starting the shell attached to the console or a serial port is very simple. The user invokes rtems_shell_init with parameters to indicate the characteristics of the task that will be executing the shell including name, stack size, and priority. The user also specifies the device that the shell is to be attached to.

This example is taken from the fileio sample test. This shell portion of this test can be run on any target which provides a console with input and output capabilities. It does not include any commands which cannot be supported on all BSPs. The source code for this test is in testsuites/samples/fileio with the shell configuration in the init.c file.

      #include <rtems/shell.h>
      void start_shell(void)
      {
        printf(" =========================\n");
        printf(" starting shell\n");
        printf(" =========================\n");
        rtems_shell_init(
          "SHLL",                       /* task name */
          RTEMS_MINIMUM_STACK_SIZE * 4, /* task stack size */
          100,                          /* task priority */
          "/dev/console",               /* device name */
          false,                        /* run forever */
          true,                         /* wait for shell to terminate */
          rtems_shell_login_check       /* login check function,
                                           use NULL to disable a login check */
        );
      }

In the above example, the call to rtems_shell_init spawns a task to run the RTEMS Shell attached to /dev/console and executing at priority 100. The caller suspends itself and lets the shell take over the console device. When the shell is exited by the user, then control returns to the caller.

1.3.2 Attached to a Socket

TBD

1.4 Functions

This section describes the Shell related C functions which are publicly available related to initialization and configuration.

1.4.1 rtems shell init - initialize the shell

CALLING SEQUENCE:

     rtems_status_code rtems_shell_init(
       const char          *task_name,
       size_t               task_stacksize,
       rtems_task_priority  task_priority,
       const char          *devname,
       bool                 forever,
       bool                 wait,
       rtems_login_check    login_check
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - Shell task spawned successfully others - to indicate a failure condition

DESCRIPTION:

This service creates a task with the specified characteristics to run the RTEMS Shell attached to the specified devname.

NOTES:

This method invokes the rtems_task_create and rtems_task_start directives and as such may return any status code that those directives may return.

2 General Commands

2.1 Introduction

The RTEMS shell has the following general commands:

• alias - Add alias for an existing command • date - Print or set current date and time • echo - Produce message in a shell script • sleep - Delay for a specified amount of time • id - show uid gid euid and egid

• tty - show ttyname

• whoami - print effective user id

• getenv - print environment variable

• setenv - set environment variable

• unsetenv - unset environment variable

• time - time command execution

• logoff - logoff from the system

• rtc - RTC driver configuration

• exit - alias for logoff command

2.2 Commands

This section details the General Commands available. A subsection is dedicated to each of the commands and describes the behavior and configuration of that command as well as providing an example usage.

2.2.1 alias - add alias for an existing command

SYNOPSYS:

     alias oldCommand newCommand

DESCRIPTION:

This command adds an alternate name for an existing command to the command set.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use alias:

     SHLL [/] $ me
     shell:me command not found
     SHLL [/] $ alias whoami me
     SHLL [/] $ me
     rtems
     SHLL [/] $ whoami
     rtems

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_ALIAS to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_ALIAS when all shell commands have been configured.

PROGRAMMING INFORMATION:

The alias is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_alias(
       int    argc,
       char **argv
     );

The configuration structure for the alias has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_ALIAS_Command;

2.2.2 date - print or set current date and time

SYNOPSYS:

     date
     date DATE TIME

DESCRIPTION:

This command operates one of two modes. When invoked with no arguments, it prints the current date and time. When invoked with both date and time arguments, it sets the current time.

The date is specified in YYYY-MM-DD format. The time is specified in HH:MM:SS format.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

This comm

EXAMPLES:

The following is an example of how to use date:

     SHLL [/] $ date
     Fri Jan  1 00:00:09 1988
     SHLL [/] $ date 2008-02-29 06:45:32
     SHLL [/] $ date
     Fri Feb 29 06:45:35 2008

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_DATE to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_DATE when all shell commands have been configured.

PROGRAMMING INFORMATION:

The date is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_date(
       int    argc,
       char **argv
     );

The configuration structure for the date has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_DATE_Command;

2.2.3 echo - produce message in a shell script

SYNOPSYS:

     echo [-n | -e] args ...

DESCRIPTION:

echo prints its arguments on the standard output, separated by spaces. Unless the -n option is present, a newline is output following the arguments. The -e option causes echo to treat the escape sequences specially, as described in the following paragraph. The -e option is the default, and is provided solely for compatibility with other systems. Only one of the options -n and -e may be given.

If any of the following sequences of characters is encountered during output, the sequence is not output. Instead, the specified action is performed:

\b A backspace character is output.

\c Subsequent output is suppressed. This is normally used at the end of the last argument to suppress the trailing newline that echo would otherwise output.

\f Output a form feed.

\n Output a newline character.

\r Output a carriage return.

\t Output a (horizontal) tab character.

\v Output a vertical tab.

\0digits Output the character whose value is given by zero to three digits. If there are zero digits, a nul character is output.

\\ Output a backslash.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

The octal character escape mechanism (\0digits) differs from the C lan- guage mechanism.

There is no way to force echo to treat its arguments literally, rather than interpreting them as options and escape sequences.

EXAMPLES:

The following is an example of how to use echo:

     SHLL [/] $ echo a b c
     a b c
     SHLL [/] $ echo

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_ECHO to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_ECHO when all shell commands have been configured.

PROGRAMMING INFORMATION:

The echo is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_echo(
       int    argc,
       char **argv
     );

The configuration structure for the echo has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_ECHO_Command;

ORIGIN:

The implementation and portions of the documentation for this command are from NetBSD 4.0.

2.2.4 sleep - delay for a specified amount of time

SYNOPSYS:

     sleep seconds
     sleep seconds nanoseconds

DESCRIPTION:

This command causes the task executing the shell to block for the specified number of

seconds and nanoseconds.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

This command is implemented using the nanosleep() method.

The command line interface is similar to the sleep command found on POSIX systems but the addition of the nanoseconds parameter allows fine grained delays in shell scripts without adding another command such as usleep.

EXAMPLES:

The following is an example of how to use sleep:

     SHLL [/] $ sleep 10
     SHLL [/] $ sleep 0 5000000

It is not clear from the above but there is a ten second pause after executing the first command before the prompt is printed. The second command completes very quickly from a human perspective and there is no noticeable delay in the prompt being printed.

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_SLEEP to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_SLEEP when all shell commands have been configured.

PROGRAMMING INFORMATION:

The sleep is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_sleep(
       int    argc,
       char **argv
     );

The configuration structure for the sleep has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_SLEEP_Command;

2.2.5 id - show uid gid euid and egid

SYNOPSYS:

     id

DESCRIPTION:

This command prints the user identity. This includes the user id (uid), group id (gid), effective user id (euid), and effective group id (egid).

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

Remember there is only one POSIX process in a single processor RTEMS application. Each thread may have its own user identity and that identity is used by the filesystem to enforce permissions.

EXAMPLES:

The first example of the id command is from a session logged in as the normal user rtems:

     SHLL [/] # id
     uid=1(rtems),gid=1(rtems),euid=1(rtems),egid=1(rtems)

The second example of the id command is from a session logged in as the root user:

     SHLL [/] # id
     uid=0(root),gid=0(root),euid=0(root),egid=0(root)

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_ID to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_ID when all shell commands have been configured.

PROGRAMMING INFORMATION:

The id is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_id(
       int    argc,
       char **argv
     );

The configuration structure for the id has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_ID_Command;

2.2.6 tty - show ttyname SYNOPSYS:

     tty

DESCRIPTION:

This command prints the file name of the device connected to standard input.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use tty:

     SHLL [/] $ tty
     /dev/console

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_TTY to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_TTY when all shell commands have been configured.

PROGRAMMING INFORMATION:

The tty is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_tty(
       int    argc,
       char **argv
     );

The configuration structure for the tty has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_TTY_Command;

2.2.7 whoami - print effective user id SYNOPSYS:

     whoami

DESCRIPTION:

This command displays the user name associated with the current effective user id.

EXIT STATUS:

This command always succeeds.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use whoami:

     SHLL [/] $ whoami
     rtems

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_WHOAMI to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_WHOAMI when all shell commands have been configured.

PROGRAMMING INFORMATION:

The whoami is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_whoami(
       int    argc,
       char **argv
     );

The configuration structure for the whoami has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_WHOAMI_Command;

2.2.8 getenv - print environment variable SYNOPSYS:

     getenv variable

DESCRIPTION:

This command is used to display the value of a variable in the set of environment variables.

EXIT STATUS:

This command will return 1 and print a diagnostic message if a failure occurs.

NOTES:

The entire RTEMS application shares a single set of environment variables.

EXAMPLES:

The following is an example of how to use getenv:

     SHLL [/] $ getenv BASEPATH
     /mnt/hda1

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_GETENV to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_GETENV when all shell commands have been configured.

PROGRAMMING INFORMATION:

The getenv is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_getenv(
       int    argc,
       char **argv
     );

The configuration structure for the getenv has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_GETENV_Command;

2.2.9 setenv - set environment variable SYNOPSYS:

     setenv variable [value]

DESCRIPTION:

This command is used to add a new variable to the set of environment variables or to modify the variable of an already existing variable. If the value is not provided, the variable will be set to the empty string.

EXIT STATUS:

This command will return 1 and print a diagnostic message if a failure occurs.

NOTES:

The entire RTEMS application shares a single set of environment variables.

EXAMPLES:

The following is an example of how to use setenv:

     SHLL [/] $ setenv BASEPATH /mnt/hda1

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_SETENV to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_SETENV when all shell commands have been configured.

PROGRAMMING INFORMATION:

The setenv is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_setenv(
       int    argc,
       char **argv
     );

The configuration structure for the setenv has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_SETENV_Command;

2.2.10 unsetenv - unset environment variable SYNOPSYS:

     unsetenv variable

DESCRIPTION:

This command is remove to a variable from the set of environment variables.

EXIT STATUS:

This command will return 1 and print a diagnostic message if a failure occurs.

NOTES:

The entire RTEMS application shares a single set of environment variables.

EXAMPLES:

The following is an example of how to use unsetenv:

     SHLL [/] $ unsetenv BASEPATH

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_UNSETENV to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_UNSETENV when all shell commands have been configured.

PROGRAMMING INFORMATION:

The unsetenv is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_unsetenv(
       int    argc,
       char **argv
     );

The configuration structure for the unsetenv has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_UNSETENV_Command;

2.2.11 time - time command execution SYNOPSYS:

     time command [argument ...]

DESCRIPTION:

The time command executes and times a command. After the command finishes, time writes the total time elapsed. Times are reported in seconds.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

None.

EXAMPLES:

The following is an example of how to use time:

     SHLL [/] $ time cp -r /nfs/directory /c

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_TIME to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_TIME when all shell commands have been configured.

PROGRAMMING INFORMATION:

The time is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_time(
       int    argc,
       char **argv
     );

The configuration structure for the time has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_TIME_Command;

2.2.12 logoff - logoff from the system

SYNOPSYS:

     logoff

DESCRIPTION:

This command logs the user out of the shell.

EXIT STATUS:

This command does not return.

NOTES:

The system behavior when the shell is exited depends upon how the shell was initiated. The typical behavior is that a login prompt will be displayed for the next login attempt or that the connection will be dropped by the RTEMS system.

EXAMPLES:

The following is an example of how to use logoff:

     SHLL [/] $ logoff
     logoff from the system...

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_LOGOFF to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_LOGOFF when all shell commands have been configured.

PROGRAMMING INFORMATION:

The logoff is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_logoff(
       int    argc,
       char **argv
     );

The configuration structure for the logoff has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_LOGOFF_Command;

2.2.13 rtc - RTC driver configuration SYNOPSYS:

     rtc

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_RTC to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_RTC when all shell commands have been configured.

2.2.14 exit - exit the shell SYNOPSYS:

     exit

DESCRIPTION:

This command causes the shell interpreter to exit.

EXIT STATUS:

This command does not return.

NOTES:

In contrast to Section 2.2.12 [General Commands logoff - logoff from the system], page 20, this command is built into the shell interpreter loop.

EXAMPLES:

The following is an example of how to use exit:

     SHLL [/] $ exit
     Shell exiting

CONFIGURATION:

This command is always present and cannot be disabled.

PROGRAMMING INFORMATION:

The exit is implemented directly in the shell interpreter. There is no C routine associated with it.

3 File and Directory Commands

3.1 Introduction

The RTEMS shell has the following file and directory commands:

• umask - Set file mode creation mask • cp - copy files

• mv - move files

• pwd - print work directory

• ls - list files in the directory

• chdir - change the current directory • mkdir - create a directory

• rmdir - remove empty directories • ln - make links

• mknod - make device special file

• chroot - change the root directory • chmod - change permissions of a file • cat - display file contents

• msdosfmt - format disk

• rm - remove files

• mount - mount disk

• unmount - unmount disk

• blksync - sync the block driver

• dd - format disks

• hexdump - format disks

• fdisk - format disks

• dir - alias for ls

• mkrfs - format RFS file system

• cd - alias for chdir

3.2 Commands

This section details the File and Directory Commands available. A subsection is dedicated to each of the commands and describes the behavior and configuration of that command as well as providing an example usage.

3.2.1 umask - set file mode creation mask

SYNOPSYS:

     umask [new_umask]

DESCRIPTION:

This command sets the user file creation mask to new_umask. The argument new_umask may be octal, hexadecimal, or decimal.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

This command does not currently support symbolic mode masks.

EXAMPLES:

The following is an example of how to use umask:

     SHLL [/] $ umask
     022
     SHLL [/] $ umask 0666
     0666
     SHLL [/] $ umask
     0666

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_UMASK to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_UMASK when all shell commands have been configured.

PROGRAMMING INFORMATION:

The umask is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_umask(
       int    argc,
       char **argv
     );

The configuration structure for the umask has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_UMASK_Command;

3.2.2 cp - copy files

SYNOPSYS:

cp [-R [-H | -L | -P]] [-f | -i] [-pv] src target
cp [-R [-H | -L] ] [-f | -i] [-NpPv] source_file ... target_directory

DESCRIPTION:

In the first synopsis form, the cp utility copies the contents of the source file to the target file. In the second synopsis form, the contents of each named source file is copied to the destination target directory. The names of the files themselves are not changed. If cp detects an attempt to copy a file to itself, the copy will fail.

The following options are available:

-f For each existing destination pathname, attempt to overwrite it. If permissions do not allow copy to succeed, remove it and create a new file, without prompting for confirmation. (The -i option is ignored if the -f option is specified.)

-H If the -R option is specified, symbolic links on the command line are followed. (Symbolic links encountered in the tree traversal are not followed.)

-i Causes cp to write a prompt to the standard error output before copying a file that would overwrite an existing file. If the response from the standard input begins with the character ’y’, the file copy is attempted.

-L If the -R option is specified, all symbolic links are followed.

-N When used with -p, do not copy file flags.

-P No symbolic links are followed.

-p Causes cp to preserve in the copy as many of the modification time, access time, file flags, file mode, user ID, and group ID as allowed by permissions. If the user ID and group ID cannot be preserved, no error message is displayed and the exit value is not altered. If the source file has its set user ID bit on and the user ID cannot be preserved, the set user ID bit is not preserved in the copy’s permissions. If the source file has its set group ID bit on and the group ID cannot be preserved, the set group ID bit is not preserved in the copy’s permissions. If the source file has both its set user ID and set group ID bits on, and either the user ID or group ID cannot be preserved, neither the set user ID or set group ID bits are preserved in the copy’s permissions.

-R If source file designates a directory, cp copies the directory and the entire subtree connected at that point. This option also causes symbolic links to be copied, rather than indirected through, and for cp to create special files rather than copying them as normal files. Created directories have the same mode as the corresponding source directory, unmodified by the process’s umask.

-v Cause cp to be verbose, showing files as they are copied.

For each destination file that already exists, its contents are overwritten if permissions allow, but its mode, user ID, and group ID are unchanged.

In the second synopsis form, target directory must exist unless there is only one named source file which is a directory and the -R flag is specified.

If the destination file does not exist, the mode of the source file is used as modified by the file mode creation mask (umask, see csh(1)). If the source file has its set user ID bit on, that bit is removed unless both the source file and the destination file are owned by the same user. If the source file has its set group ID bit on, that bit is removed unless both the source file and the destination file are in the same group and the user is a member of that group. If both the set user ID and set group ID bits are set, all of the above conditions must be fulfilled or both bits are removed.

Appropriate permissions are required for file creation or overwriting.

Symbolic links are always followed unless the -R flag is set, in which case symbolic links are not followed, by default. The -H or -L flags (in conjunction with the -R flag), as well as the -P flag cause symbolic links to be followed as described above. The -H and -L options are ignored unless the -R option is specified. In addition, these options override eachsubhedading other and the command’s actions are determined by the last one specified.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use cp to copy a file to a new name in the current directory:

     SHLL [/] # cat joel
     cat: joel: No such file or directory
     SHLL [/] # cp etc/passwd joel
     SHLL [/] # cat joel
     root:*:0:0:root::/:/bin/sh
     rtems:*:1:1:RTEMS Application::/:/bin/sh
     tty:!:2:2:tty owner::/:/bin/false
     SHLL [/] # ls
     drwxr-xr-x   1   root   root         536 Jan 01 00:00 dev/
     drwxr-xr-x   1   root   root        1072 Jan 01 00:00 etc/
     -rw-r--r--   1   root   root         102 Jan 01 00:00 joel
     3 files 1710 bytes occupied

The following is an example of how to use cp to copy one or more files to a destination directory and use the same basename in the destination directory:

SHLL [/] # mkdir tmp
SHLL [/] # ls tmp
0 files 0 bytes occupied
SHLL [/] # cp /etc/passwd tmp
SHLL [/] # ls /tmp
-rw-r--r--   1   root   root         102 Jan 01 00:01 passwd
1 files 102 bytes occupied
SHLL [/] # cp /etc/passwd /etc/group /tmp
SHLL [/] # ls /tmp
-rw-r--r--   1   root   root         102 Jan 01 00:01 passwd
-rw-r--r--   1   root   root          42 Jan 01 00:01 group
2 files 144 bytes occupied
SHLL [/] #

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_CP to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_CP when all shell commands have been configured.

PROGRAMMING INFORMATION:

The cp command is implemented by a C language function which has the following prototype:

int rtems_shell_rtems_main_cp(
  int    argc,
  char **argv
);

The configuration structure for the cp has the following prototype:

extern rtems_shell_cmd_t rtems_shell_CP_Command;

ORIGIN:

The implementation and portions of the documentation for this command are from NetBSD 4.0.

3.2.3 mv - move files

SYNOPSYS:

     mv [-fiv] source_file target_file
     mv [-fiv] source_file... target_file

DESCRIPTION:

In its first form, the mv utility renames the file named by the source operand to the destination path named by the target operand. This form is assumed when the last operand does not name an already existing directory.

In its second form, mv moves each file named by a source operand to a destination file in the existing directory named by the directory operand. The destination path for each operand is the pathname produced by the concatenation of the last operand, a slash, and the final pathname component of the named file.

The following options are available:

-f Do not prompt for confirmation before overwriting the destination path.

-i Causes mv to write a prompt to standard error before moving a file that would overwrite an existing file. If the response from the standard input begins with the character ’y’, the move is attempted.

-v Cause mv to be verbose, showing files as they are processed.

The last of any -f or -i options is the one which affects mv’s behavior.

It is an error for any of the source operands to specify a nonexistent file or directory.

It is an error for the source operand to specify a directory if the target exists and is not a directory.

If the destination path does not have a mode which permits writing, mv prompts the user for confirmation as specified for the -i option.

Should the rename call fail because source and target are on different file systems, mv will remove the destination file, copy the source file to the destination, and then remove the source. The effect is roughly equivalent to:

     rm -f destination_path && \
     cp -PRp source_file destination_path && \
     rm -rf source_file

EXIT STATUS:

The mv utility exits 0 on success, and >0 if an error occurs.

NOTES:

NONE

EXAMPLES:

     SHLL [/] mv /dev/console /dev/con1

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_MV to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_MV when all shell commands have been configured.

PROGRAMMING INFORMATION:

The mv command is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_mv(
       int    argc,
       char **argv
     );

The configuration structure for the mv has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_MV_Command;

ORIGIN:

The implementation and portions of the documentation for this command are from NetBSD 4.0.

3.2.4 pwd - print work directory

SYNOPSYS:

     pwd

DESCRIPTION:

This command prints the fully qualified filename of the current working directory.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use pwd:

     SHLL [/] $ pwd
     /
     SHLL [/] $ cd dev
     SHLL [/dev] $ pwd
     /dev

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_PWD to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_PWD when all shell commands have been configured.

PROGRAMMING INFORMATION:

The pwd is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_pwd(
       int    argc,
       char **argv
     );

The configuration structure for the pwd has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_PWD_Command;

3.2.5 ls - list files in the directory

SYNOPSYS:

ls [dir]

DESCRIPTION:

This command displays the contents of the specified directory. If no arguments are given, then it displays the contents of the current working directory.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

This command currently does not display information on a set of files like the POSIX ls(1). It only displays the contents of entire directories.

EXAMPLES:

The following is an example of how to use ls:

SHLL [/] $ ls
drwxr-xr-x   1   root   root         536 Jan 01 00:00 dev/
drwxr-xr-x   1   root   root        1072 Jan 01 00:00 etc/
2 files 1608 bytes occupied
SHLL [/] $ ls etc
-rw-r--r--   1   root   root         102 Jan 01 00:00 passwd
-rw-r--r--   1   root   root          42 Jan 01 00:00 group
-rw-r--r--   1   root   root          30 Jan 01 00:00 issue
-rw-r--r--   1   root   root          28 Jan 01 00:00 issue.net
4 files 202 bytes occupied
SHLL [/] $ ls dev etc
-rwxr-xr-x   1  rtems   root           0 Jan 01 00:00 console
-rwxr-xr-x   1   root   root           0 Jan 01 00:00 console_b

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_LS to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_LS when all shell commands have been configured.

PROGRAMMING INFORMATION:

The ls is implemented by a C language function which has the following prototype:

int rtems_shell_rtems_main_ls(
  int    argc,
  char **argv
);

The configuration structure for the ls has the following prototype:

32 RTEMS Shell User’s Guide

     extern rtems_shell_cmd_t rtems_shell_LS_Command;

3.2.6 chdir - change the current directory

SYNOPSYS:

     chdir [dir]

DESCRIPTION:

This command is used to change the current working directory to the specified directory. If no arguments are given, the current working directory will be changed to /.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use chdir:

     SHLL [/] $ pwd
     /
     SHLL [/] $ chdir etc
     SHLL [/etc] $ pwd
     /etc

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_CHDIR to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_CHDIR when all shell commands have been configured.

PROGRAMMING INFORMATION:

The chdir is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_chdir(
       int    argc,
       char **argv
     );

The configuration structure for the chdir has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_CHDIR_Command;

3.2.7 mkdir - create a directory

SYNOPSYS:

     mkdir  dir [dir1 .. dirN]

DESCRIPTION:

This command creates the set of directories in the order they are specified on the command line. If an error is encountered making one of the directories, the command will continue to attempt to create the remaining directories on the command line.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

If this command is invoked with no arguments, nothing occurs.

The user must have sufficient permissions to create the directory. For the fileio test provided with RTEMS, this means the user must login as root not rtems.

EXAMPLES:

The following is an example of how to use mkdir:

     SHLL [/] # ls
     drwxr-xr-x   1   root   root         536 Jan 01 00:00 dev/
     drwxr-xr-x   1   root   root        1072 Jan 01 00:00 etc/
     2 files 1608 bytes occupied
     SHLL [/] # mkdir joel
     SHLL [/] # ls joel
     0 files 0 bytes occupied
     SHLL [/] # cp etc/passwd joel
     SHLL [/] # ls joel
     -rw-r--r--   1   root   root         102 Jan 01 00:02 passwd
     1 files 102 bytes occupied

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_MKDIR to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_MKDIR when all shell commands have been configured.

PROGRAMMING INFORMATION:

The mkdir is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_mkdir(
       int    argc,
       char **argv
     );

Chapter 3: File and Directory Commands 35 The configuration structure for the mkdir has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_MKDIR_Command;

3.2.8 rmdir - remove empty directories

SYNOPSYS:

     rmdir  [dir1 .. dirN]

DESCRIPTION:

This command removes the specified set of directories. If no directories are provided on the command line, no actions are taken.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

This command is a implemented using the rmdir(2) system call and all reasons that call may fail apply to this command.

EXAMPLES:

The following is an example of how to use rmdir:

     SHLL [/] # mkdir joeldir
     SHLL [/] # rmdir joeldir
     SHLL [/] # ls joeldir
     joeldir: No such file or directory.

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_RMDIR to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_RMDIR when all shell commands have been configured.

PROGRAMMING INFORMATION:

The rmdir is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_rmdir(
       int    argc,
       char **argv
     );

The configuration structure for the rmdir has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_RMDIR_Command;

3.2.9 ln - make links

SYNOPSYS:

     ln [-fhinsv] source_file [target_file]
     ln [-fhinsv] source_file ... target_dir

DESCRIPTION:

The ln utility creates a new directory entry (linked file) which has the same modes as the original file. It is useful for maintaining multiple copies of a file in many places at once without using up storage for the “copies”; instead, a link “points” to the original copy. There are two types of links; hard links and symbolic links. How a link “points” to a file is one of the differences between a hard or symbolic link.

The options are as follows:

-f Unlink any already existing file, permitting the link to occur.

-h If the target file or target dir is a symbolic link, do not follow it. This is most useful with the -f option, to replace a symlink which may point to a directory.

-i Cause ln to write a prompt to standard error if the target file exists. If the response from the standard input begins with the character ‘y’ or ‘Y’, then unlink the target file so that the link may occur. Otherwise, do not attempt the link. (The -i option overrides any previous -f options.)

-n Same as -h, for compatibility with other ln implementations.

-s Create a symbolic link.

-v Cause ln to be verbose, showing files as they are processed.

By default ln makes hard links. A hard link to a file is indistinguishable from the original directory entry; any changes to a file are effective independent of the name used to reference the file. Hard links may not normally refer to directories and may not span file systems.

A symbolic link contains the name of the file to which it is linked. The referenced file is used when an open operation is performed on the link. A stat on a symbolic link will return the linked-to file; an lstat must be done to obtain information about the link. The readlink call may be used to read the contents of a symbolic link. Symbolic links may span file systems and may refer to directories.

Given one or two arguments, ln creates a link to an existing file source file. If target file is given, the link has that name; target file may also be a directory in which to place the link; otherwise it is placed in the current directory. If only the directory is specified, the link will be made to the last component of source file.

Given more than two arguments, ln makes links in target dir to all the named source files. The links made will have the same name as the files being linked to.

EXIT STATUS:

The ln utility exits 0 on success, and >0 if an error occurs.

NOTES:

NONE

EXAMPLES:

     SHLL [/] ln -s /dev/console /dev/con1

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_LN to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_LN when all shell commands have been configured.

PROGRAMMING INFORMATION:

The ln command is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_ln(
       int    argc,
       char **argv
     );

The configuration structure for the ln has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_LN_Command;

ORIGIN:

The implementation and portions of the documentation for this command are from NetBSD 4.0.

3.2.10 mknod - make device special file

SYNOPSYS:

mknod [-rR] [-F fmt] [-g gid] [-m mode] [-u uid] name [c | b]
[driver | major] minor
mknod [-rR] [-F fmt] [-g gid] [-m mode] [-u uid] name [c | b]
major unit subunit
mknod [-rR] [-g gid] [-m mode] [-u uid] name [c | b] number
mknod [-rR] [-g gid] [-m mode] [-u uid] name p

DESCRIPTION:

The mknod command creates device special files, or fifos. Normally the shell script /dev/MAKEDEV is used to create special files for commonly known devices; it executes mknod with the appropriate arguments and can make all the files required for the device.

To make nodes manually, the arguments are:

-r Replace an existing file if its type is incorrect.

-R Replace an existing file if its type is incorrect. Correct the mode, user and group.

-g gid Specify the group for the device node. The gid operand may be a numeric group ID or a group name. If a group name is also a numeric group ID, the operand is used as a group name. Precede a numeric group ID with a # to stop it being treated as a name.

-m mode Specify the mode for the device node. The mode may be absolute or symbolic, see chmod.

-u uid Specify the user for the device node. The uid operand may be a numeric user ID or a user name. If a user name is also a numeric user ID, the operand is used as a user name. Precede a numeric user ID with a # to stop it being treated as a name.

name Device name, for example “tty” for a termios serial device or “hd” for a disk.

b | c | p Type of device. If the device is a block type device such as a tape or disk drive which needs both cooked and raw special files, the type is b. All other devices are character type devices, such as terminal and pseudo devices, and are type c. Specifying p creates fifo files.

driver | major The major device number is an integer number which tells the kernel which device driver entry point to use. If the device driver is configured into the current kernel it may be specified by driver name or major number.

minor The minor device number tells the kernel which one of several similar devices the node corresponds to; for example, it may be a specific serial port or pty. unit and subunit The unit and subunit numbers select a subset of a device; for example,

the unit may specify a particular disk, and the subunit a partition on that disk. (Currently this form of specification is only supported by the bsdos format, for compatibility with the BSD/OS mknod).

number A single opaque device number. Useful for netbooted computers which require device numbers packed in a format that isn’t supported by -F.

EXIT STATUS:

The mknod utility exits 0 on success, and >0 if an error occurs.

NOTES:

NONE

EXAMPLES:

     SHLL [/] mknod c 3 0 /dev/ttyS10

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_MKNOD to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_MKNOD when all shell commands have been configured.

PROGRAMMING INFORMATION:

The mknod command is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_mknod(
       int    argc,
       char **argv
     );

The configuration structure for the mknod has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_MKNOD_Command;

ORIGIN:

The implementation and portions of the documentation for this command are from NetBSD 4.0.

3.2.11 chroot - change the root directory

SYNOPSYS:

     chroot [dir]

DESCRIPTION:

This command changes the root directory to dir for subsequent commands.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

The destination directory dir must exist.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use chroot and the impact it has on the environment for subsequent command invocations:

     SHLL [/] $ cat passwd
     cat: passwd: No such file or directory
     SHLL [/] $ chroot etc
     SHLL [/] $ cat passwd
     root:*:0:0:root::/:/bin/sh
     rtems:*:1:1:RTEMS Application::/:/bin/sh
     tty:!:2:2:tty owner::/:/bin/false
     SHLL [/] $ cat /etc/passwd
     cat: /etc/passwd: No such file or directory

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_CHROOT to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_CHROOT when all shell commands have been configured.

PROGRAMMING INFORMATION:

The chroot is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_chroot(
       int    argc,
       char **argv
     );

The configuration structure for the chroot has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_CHROOT_Command;

3.2.12 chmod - change permissions of a file

SYNOPSYS:

chmod permissions file1 [file2...]

DESCRIPTION:

This command changes the permissions on the files specified to the indicated permissions. The permission values are POSIX based with owner, group, and world having individual read, write, and executive permission bits.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

The chmod command only takes numeric representations of the permissions.

EXAMPLES:

The following is an example of how to use chmod:

SHLL [/] # cd etc
SHLL [/etc] # ls
-rw-r--r--   1   root   root         102 Jan 01 00:00 passwd
-rw-r--r--   1   root   root          42 Jan 01 00:00 group
-rw-r--r--   1   root   root          30 Jan 01 00:00 issue
-rw-r--r--   1   root   root          28 Jan 01 00:00 issue.net
4 files 202 bytes occupied
SHLL [/etc] # chmod 0777 passwd
SHLL [/etc] # ls
-rwxrwxrwx   1   root   root         102 Jan 01 00:00 passwd
-rw-r--r--   1   root   root          42 Jan 01 00:00 group
-rw-r--r--   1   root   root          30 Jan 01 00:00 issue
-rw-r--r--   1   root   root          28 Jan 01 00:00 issue.net
4 files 202 bytes occupied
SHLL [/etc] # chmod 0322 passwd
SHLL [/etc] # ls
--wx-w--w-   1 nouser   root         102 Jan 01 00:00 passwd
-rw-r--r--   1 nouser   root          42 Jan 01 00:00 group
-rw-r--r--   1 nouser   root          30 Jan 01 00:00 issue
-rw-r--r--   1 nouser   root          28 Jan 01 00:00 issue.net
4 files 202 bytes occupied
SHLL [/etc] # chmod 0644 passwd
SHLL [/etc] # ls
-rw-r--r--   1   root   root         102 Jan 01 00:00 passwd
-rw-r--r--   1   root   root          42 Jan 01 00:00 group
-rw-r--r--   1   root   root          30 Jan 01 00:00 issue
-rw-r--r--   1   root   root          28 Jan 01 00:00 issue.net
4 files 202 bytes occupied

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_CHMOD to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_CHMOD when all shell commands have been configured.

PROGRAMMING INFORMATION:

The chmod is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_chmod(
       int    argc,
       char **argv
     );

The configuration structure for the chmod has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_CHMOD_Command;

3.2.13 cat - display file contents

SYNOPSYS:

     cat file1 [file2 .. fileN]

DESCRIPTION:

This command displays the contents of the specified files.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

It is possible to read the input from a device file using cat.

EXAMPLES:

The following is an example of how to use cat:

     SHLL [/] # cat /etc/passwd
     root:*:0:0:root::/:/bin/sh
     rtems:*:1:1:RTEMS Application::/:/bin/sh
     tty:!:2:2:tty owner::/:/bin/false

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_CAT to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_CAT when all shell commands have been configured.

PROGRAMMING INFORMATION:

The cat is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_cat(
       int    argc,
       char **argv
     );

The configuration structure for the cat has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_CAT_Command;

3.2.14 rm - remove files

SYNOPSYS:

     rm file1 [file2 ... fileN]

DESCRIPTION:

This command deletes a name from the filesystem. If the specified file name was the last link to a file and there are no open file descriptor references to that file, then it is deleted and the associated space in the file system is made available for subsequent use.

If the filename specified was the last link to a file but there are open file descriptor references to it, then the file will remain in existence until the last file descriptor referencing it is closed.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use rm:

     SHLL [/] # cp /etc/passwd tmpfile
     SHLL [/] # cat tmpfile
     root:*:0:0:root::/:/bin/sh
     rtems:*:1:1:RTEMS Application::/:/bin/sh
     tty:!:2:2:tty owner::/:/bin/false
     SHLL [/] # rm tmpfile
     SHLL [/] # cat tmpfile
     cat: tmpfile: No such file or directory

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_RM to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_RM when all shell commands have been configured.

PROGRAMMING INFORMATION:

The rm is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_rm(
       int    argc,
       char **argv
     );

The configuration structure for the rm has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_RM_Command;

3.2.15 mount - mount disk SYNOPSYS:

     mount [-t fstype] [-r] [-L] device path

DESCRIPTION:

The mount command will mount a block device to a mount point using the specified file system. The files systems are:

• msdos - MSDOS File System

• tftp - TFTP Network File System

• ftp - FTP Network File System

• nfs - Network File System

• rfs - RTEMS File System When the file system type is ’msdos’ or ’rfs’ the driver is a "block device driver" node present in the file system. The driver is ignored with the ’tftp’ and ’ftp’ file systems. For the ’nfs’ file system the driver is the ’host:/path’ string that described NFS host and the exported file system path.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

The mount point must exist. The services offered by each file-system vary. For example you cannot list the directory of a TFTP file-system as this server is not provided in the TFTP protocol. You need to check each file-system’s documentation for the services provided.

EXAMPLES:

Mount the Flash Disk driver to the ’/fd’ mount point:

     SHLL [/] $ mount -t msdos /dev/flashdisk0 /fd

Mount the NFS file system exported path ’bar’ by host ’foo’:

     $ mount -t nfs foo:/bar /nfs

Mount the TFTP file system on ’/tftp’:

     $ mount -t tftp /tftp

To access the TFTP files on server ’10.10.10.10’:

     $ cat /tftp/10.10.10.10/test.txt

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_MOUNT to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_MOUNT when all shell commands have been configured.

The mount command includes references to file-system code. If you do not wish to include file-system that you do not use do not define the mount command support for that filesystem. The file-system mount command defines are:

• msdos - CONFIGURE SHELL MOUNT MSDOS

• tftp - CONFIGURE SHELL MOUNT TFTP

• ftp - CONFIGURE SHELL MOUNT FTP

• nfs - CONFIGURE SHELL MOUNT NFS

• rfs - CONFIGURE SHELL MOUNT RFS

An example configuration is:

     #define CONFIGURE_SHELL_MOUNT_MSDOS
     #ifdef RTEMS_NETWORKING
       #define CONFIGURE_SHELL_MOUNT_TFTP
       #define CONFIGURE_SHELL_MOUNT_FTP
       #define CONFIGURE_SHELL_MOUNT_NFS
       #define CONFIGURE_SHELL_MOUNT_RFS
     #endif

PROGRAMMING INFORMATION:

The mount is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_mount(
       int    argc,
       char **argv
     );

The configuration structure for the mount has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_MOUNT_Command;

3.2.16 unmount - unmount disk SYNOPSYS:

     unmount path

DESCRIPTION:

This command unmounts the device at the specified path.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

TBD - Surely there must be some warnings to go here.

EXAMPLES:

The following is an example of how to use unmount:

     EXAMPLE_TBD

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_UNMOUNT to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_UNMOUNT when all shell commands have been configured.

PROGRAMMING INFORMATION:

The unmount is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_unmount(
       int    argc,
       char **argv
     );

The configuration structure for the unmount has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_UNMOUNT_Command;

3.2.17 blksync - sync the block driver SYNOPSYS:

     blksync driver

DESCRIPTION:

This command XXX

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use blksync:

     EXAMPLE_TBD

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_BLKSYNC to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_BLKSYNC when all shell commands have been configured.

PROGRAMMING INFORMATION:

The blksync is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_blksync(
       int    argc,
       char **argv
     );

The configuration structure for the blksync has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_BLKSYNC_Command;

3.2.18 dd - convert and copy a file

SYNOPSYS:

dd [operands ...]

DESCRIPTION:

The dd utility copies the standard input to the standard output. Input data is read and written in 512-byte blocks. If input reads are short, input from multiple reads are aggregated to form the output block. When finished, dd displays the number of complete and partial input and output blocks and truncated input records to the standard error output.

The following operands are available:

bs=n Set both input and output block size, superseding the ibs and obs operands. If no conversion values other than noerror, notrunc or sync are specified, then each input block is copied to the output as a single block without any aggregation of short blocks.

cbs=n Set the conversion record size to n bytes. The conversion record size is required by the record oriented conversion values.

count=n Copy only n input blocks.

files=n Copy n input files before terminating. This operand is only applicable when the input device is a tape.

ibs=n Set the input block size to n bytes instead of the default 512.

if=file Read input from file instead of the standard input.

obs=n Set the output block size to n bytes instead of the default 512.

of=file Write output to file instead of the standard output. Any regular output file is truncated unless the notrunc conversion value is specified. If an initial portion of the output file is skipped (see the seek operand) the output file is truncated at that point.

seek=n Seek n blocks from the beginning of the output before copying. On non-tape devices, a lseek operation is used. Otherwise, existing blocks are read and the data discarded. If the seek operation is past the end of file, space from the current end of file to the specified offset is filled with blocks of NUL bytes.

skip=n Skip n blocks from the beginning of the input before copying. On input which supports seeks, a lseek operation is used. Otherwise, input data is read and discarded. For pipes, the correct number of bytes is read. For all other devices, the correct number of blocks is read without distinguishing between a partial or complete block being read.

progress=n Switch on display of progress if n is set to any non-zero value. This will cause a “.” to be printed (to the standard error output) for every n full or partial blocks written to the output file. conv=value[,value...] Where value is one of the symbols from the following list.

ascii, oldascii The same as the unblock value except that characters are translated from EBCDIC to ASCII before the records are converted. (These values imply unblock if the operand cbs is also specified.) There are two conversion maps for ASCII. The value ascii specifies the recom- mended one which is compatible with AT&T System V UNIX. The value oldascii specifies the one used in historic AT&T and pre 4.3BSD-Reno systems.

block Treats the input as a sequence of newline or end-of-file terminated variable length records independent of input and output block boundaries. Any trailing newline character is discarded. Each input record is converted to a fixed length output record where the length is specified by the cbs operand. Input records shorter than the conversion record size are padded with spaces. Input records longer than the conversion record size are truncated. The number of truncated input records, if any, are reported to the standard error output at the completion of the copy.

ebcdic, ibm, oldebcdic, oldibm The same as the block value except that characters are translated from ASCII to EBCDIC after the records are converted. (These values imply block if the operand cbs is also specified.) There are four conversion maps for EBCDIC. The value ebcdic specifies the recommended one which is compatible with AT&T System V UNIX. The value ibm is a slightly different mapping, which is compatible with the AT&T System V UNIX ibm value. The values oldebcdic and oldibm are maps used in historic AT&T and pre 4.3BSD-Reno systems.

lcase Transform uppercase characters into lowercase characters.

noerror Do not stop processing on an input error. When an input error occurs, a diagnostic message followed by the current input and output block counts will be written to the standard error output in the same format as the standard completion message. If the sync conversion is also specified, any missing input data will be replaced with NUL bytes (or with spaces if a block oriented conversion value was specified) and processed as a normal input buffer. If the sync conversion is not specified, the input block is omitted from the output. On input files which are not tapes or pipes, the file offset will be positioned past the block in which the error occurred using lseek(2).

notrunc Do not truncate the output file. This will preserve any blocks in the output file not explicitly written by dd. The notrunc value is not supported for tapes.

osync Pad the final output block to the full output block size. If the input file is not a multiple of the output block size after conversion, this conversion forces the final output block to be the same size as preceding blocks for use on devices that require regularly sized blocks to be written. This option is incompatible with use of the bs=n block size specification.

sparse If one or more non-final output blocks would consist solely of NUL bytes, try to seek the output file by the required space instead of filling them with NULs. This results in a sparse file on some file systems.

swab Swap every pair of input bytes. If an input buffer has an odd number of bytes, the last byte will be ignored during swapping.

sync Pad every input block to the input buffer size. Spaces are used for pad bytes if a block oriented conversion value is specified, otherwise NUL bytes are used.

ucase Transform lowercase characters into uppercase characters.

unblock Treats the input as a sequence of fixed length records independent of input and output block boundaries. The length of the input records is specified by the cbs operand. Any trailing space characters are discarded and a newline character is appended.

Where sizes are specified, a decimal number of bytes is expected. Two or more numbers may be separated by an “x” to indicate a product. Each number may have one of the following optional suffixes:

b Block; multiply by 512

k Kibi; multiply by 1024 (1 KiB)

m Mebi; multiply by 1048576 (1 MiB)

g Gibi; multiply by 1073741824 (1 GiB)

t Tebi; multiply by 1099511627776 (1 TiB)

w Word; multiply by the number of bytes in an integer

When finished, dd displays the number of complete and partial input and output blocks, truncated input records and odd-length byte-swapping ritten. Partial output blocks to tape devices are considered fatal errors. Otherwise, the rest of the block will be written. Partial output blocks to character devices will produce a warning message. A truncated input block is one where a variable length record oriented conversion value was specified and the input line was too long to fit in the conversion record or was not newline terminated.

Normally, data resulting from input or conversion or both are aggregated into output blocks of the specified size. After the end of input is reached, any remaining output is written as a block. This means that the final output block may be shorter than the output block size.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use dd:

     SHLL [/] $ dd if=/nfs/boot-image of=/dev/hda1

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_DD to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_DD when all shell commands have been configured.

PROGRAMMING INFORMATION:

The dd command is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_dd(
       int    argc,
       char **argv

     );

The configuration structure for the dd has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_DD_Command;

3.2.19 hexdump - ascii/dec/hex/octal dump

SYNOPSYS:

hexdump [-bcCdovx] [-e format_string] [-f format_file] [-n length]
[-s skip] file ...

DESCRIPTION:

The hexdump utility is a filter which displays the specified files, or the standard input, if no files are specified, in a user specified format.

The options are as follows:

-b One-byte octal display. Display the input offset in hexadecimal, followed by sixteen space-separated, three column, zero-filled, bytes of input data, in octal, per line.

-c One-byte character display. Display the input offset in hexadecimal, followed by sixteen space-separated, three column, space-filled, characters of input data per line.

-C Canonical hex+ASCII display. Display the input offset in hexadecimal, followed by sixteen space-separated, two column, hexadecimal bytes, followed by the same sixteen bytes in % p format enclosed in “|” characters.

-d Two-byte decimal display. Display the input offset in hexadecimal, followed by eight space-separated, five column, zero-filled, two-byte units of input data, in unsigned decimal, per line.

-e format string Specify a format string to be used for displaying data.

-f format file Specify a file that contains one or more newline separated format strings. Empty lines and lines whose first non-blank character is a hash mark (#) are ignored.

-n length Interpret only length bytes of input.

-o Two-byte octal display. Display the input offset in hexadecimal, followed by eight space-separated, six column, zerofilled, two byte quantities of input data, in octal, per line.

-s offset Skip offset bytes from the beginning of the input. By default, offset is interpreted as a decimal number. With a leading 0x or 0X, offset is interpreted as a hexadecimal number, otherwise, with a leading 0, offset is interpreted as an octal number. Appending the character b, k, or m to offset causes it to be interpreted as a multiple of 512, 1024, or 1048576, respectively.

-v The -v option causes hexdump to display all input data. Without the -v option, any number of groups of output lines, which would be identical to the immediately preceding group of output lines (except for the input offsets), are replaced with a line containing a single asterisk.

-x Two-byte hexadecimal display. Display the input offset in hexadecimal, followed by eight, space separated, four column, zero-filled, two-byte quantities of input data, in hexadecimal, per line.

For each input file, hexdump sequentially copies the input to standard output, transforming the data according to the format strings specified by the -e and -f options, in the order that they were specified.

Formats

A format string contains any number of format units, separated by whitespace. A format unit contains up to three items: an iteration count, a byte count, and a format.

The iteration count is an optional positive integer, which defaults to one. Each format is applied iteration count times.

The byte count is an optional positive integer. If specified it defines the number of bytes to be interpreted by each iteration of the format.

If an iteration count and/or a byte count is specified, a single slash must be placed after the iteration count and/or before the byte count to disambiguate them. Any whitespace before or after the slash is ignored.

The format is required and must be surrounded by double quote (“ “) marks. It is interpreted as a fprintf-style format string (see fprintf ), with the following exceptions:

• An asterisk (*) may not be used as a field width or precision.

• A byte count or field precision is required for each “s” con- version character (unlike the fprintf(3) default which prints the entire string if the precision is unspecified).

• The conversion characters “h”, “l”, “n”, “p” and “q” are not supported.

• The single character escape sequences described in the C standard are supported: NUL \0 <alert character> \a <backspace> \b <form-feed> \f <newline> \n <carriage return> \r <tab> \t <vertical tab> \v

Hexdump also supports the following additional conversion strings:

a[dox] Display the input offset, cumulative across input files, of the next byte to be displayed. The appended characters d, o, and x specify the display base as decimal, octal or hexadecimal respectively.

A[dox] Identical to the a conversion string except that it is only performed once, when all of the input data has been processed.

c Output characters in the default character set. Nonprinting characters are displayed in three character, zero-padded octal, except for those representable by standard escape notation (see above), which are displayed as two character strings.

p Output characters in the default character set. Nonprinting characters are displayed as a single “.”.

u Output US ASCII characters, with the exception that control characters are displayed using the following, lower-case, names. Characters greater than 0xff, hexadecimal, are displayed as hexadecimal strings.

000 nul 001 soh 002 stx 003 etx 004 eot 005 enq 006 ack 007 bel 008 bs 009 ht 00A lf 00B vt 00C ff 00D cr 00E so 00F si 010 dle 011 dc1 012 dc2 013 dc3 014 dc4 015 nak 016 syn 017 etb 018 can 019 em 01A sub 01B esc 01C fs 01D gs 01E rs 01F us 07F del

The default and supported byte counts for the conversion characters are as follows:

% c, % p, % u, %c One byte counts only.

%d, %i, %o, %u, %X, %x Four byte default, one, two, four and eight byte counts supported.

%E, %e, %f, %G, %g Eight byte default, four byte counts supported.

The amount of data interpreted by each format string is the sum of the data required by each format unit, which is the iteration count times the byte count, or the iteration count times the number of bytes required by the format if the byte count is not specified.

The input is manipulated in “blocks”, where a block is defined as the largest amount of data specified by any format string. Format strings interpreting less than an input block’s worth of data, whose last format unit both interprets some number of bytes and does not have a specified iteration count, have the iteration count incremented until the entire input block has been processed or there is not enough data remaining in the block to satisfy the format string.

If, either as a result of user specification or hexdump modifying the iteration count as described above, an iteration count is greater than one, no trailing whitespace characters are output during the last iteration.

It is an error to specify a byte count as well as multiple conversion characters or strings unless all but one of the conversion characters or strings is a or A.

If, as a result of the specification of the -n option or end-of-file being reached, input data only partially satisfies a format string, the input block is zero-padded sufficiently to display all available data (i.e. any format units overlapping the end of data will display some number of the zero bytes).

Further output by such format strings is replaced by an equivalent number of spaces. An equivalent number of spaces is defined as the number of spaces output by an s conversion character with the same field width and precision as the original conversion character or conversion string but with any “+”, “ ”, “#” conversion flag characters removed, and referencing a NULL string.

If no format strings are specified, the default display is equivalent to specifying the -x option.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use hexdump:

     SHLL [/] $ hexdump -C -n 512 /dev/hda1

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_HEXDUMP to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_HEXDUMP when all shell commands have been configured.

PROGRAMMING INFORMATION:

The hexdump command is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_hexdump(
       int    argc,
       char **argv
     );

The configuration structure for the hexdump has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_HEXDUMP_Command;

3.2.20 fdisk - format disk SYNOPSYS:

     fdisk

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_FDISK to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_FDISK when all shell commands have been configured.

3.2.21 dir - alias for ls

SYNOPSYS:

dir [dir]

DESCRIPTION:

This command is an alias or alternate name for the ls. See Section 3.2.5 [File and Directory Commands ls - list files in the directory], page 31 for more information.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use dir:

SHLL [/] $ dir
drwxr-xr-x   1   root   root         536 Jan 01 00:00 dev/
drwxr-xr-x   1   root   root        1072 Jan 01 00:00 etc/
2 files 1608 bytes occupied
SHLL [/] $ dir etc
-rw-r--r--   1   root   root         102 Jan 01 00:00 passwd
-rw-r--r--   1   root   root          42 Jan 01 00:00 group
-rw-r--r--   1   root   root          30 Jan 01 00:00 issue
-rw-r--r--   1   root   root          28 Jan 01 00:00 issue.net
4 files 202 bytes occupied

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_DIR to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_DIR when all shell commands have been configured.

PROGRAMMING INFORMATION:

The dir is implemented by a C language function which has the following prototype:

int rtems_shell_rtems_main_dir(
  int    argc,
  char **argv
);

The configuration structure for the dir has the following prototype:

extern rtems_shell_cmd_t rtems_shell_DIR_Command;

3.2.22 mkrfs - format RFS file system

SYNOPSYS:

     mkrfs [-vsbiIo] device

DESCRIPTION:

Format the block device with the RTEMS File System (RFS). The default configuration with not parameters selects a suitable block size based on the size of the media being formatted.

The media is broken up into groups of blocks. The number of blocks in a group is based on the number of bits a block contains. The large a block the more blocks a group contains and the fewer groups in the file system.

The following options are provided:

-v Display configuration and progress of the format.

-s Set the block size in bytes.

-b The number of blocks in a group. The block count must be equal or less than the number of bits in a block.

-i Number of inodes in a group. The inode count must be equal or less than the number of bits in a block.

-I Initialise the inodes. The default is not to initialise the inodes and to rely on the inode being initialised when allocated. Initialising the inode table helps recovery if a problem appears.

-o Integer percentage of the media used by inodes. The default is 1%.

device Path of the device to format.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use mkrfs:

     SHLL [/] $ mkrfs /dev/fdda

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_MKRFS to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_MKRFS when all shell commands have been configured.

PROGRAMMING INFORMATION:

The mkrfs command is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_mkrfs(
       int    argc,
       char **argv
     );

The configuration structure for mkrfs has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_MKRFS_Command;

3.2.23 debugrfs - debug RFS file system SYNOPSYS:

     debugrfs [-hl] path command [options]

DESCRIPTION:

The command provides debugging information for the RFS file system. The options are: -h Print a help message. -l List the commands. path Path to the mounted RFS file system. The file system has to be

mounted to view to use this command.

The commands are: block start [end] Display the contents of the blocks from start to end. data Display the file system data and configuration. dir bno Process the block as a directory displaying the entries. group start [end] Display the group data from the start group to the end group. inode [-aef] [start] [end]

Display the inodes between start and end. If no start and end is provides all inodes are displayed. -a Display all inodes. That is allocated and un-

allocated inodes.

-e Search and display on inodes that have an er-

ror.

-f Force display of inodes, even when in error.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use debugrfs:

     SHLL [/] $ debugrfs /c data

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_DEBUGRFS to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_DEBUGRFS when all shell commands have been configured.

PROGRAMMING INFORMATION:

The debugrfs command is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_debugrfs(
       int    argc,
       char **argv
     );

The configuration structure for debugrfs has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_DEBUGRFS_Command;

3.2.24 cd - alias for chdir

SYNOPSYS:

cd directory

DESCRIPTION:

This command is an alias or alternate name for the chdir. See Section 3.2.6 [File and Directory Commands chdir - change the current directory], page 33 for more information.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use cd:

SHLL [/] $ cd etc
SHLL [/etc] $ cd /
SHLL [/] $ cd /etc
SHLL [/etc] $ pwd
/etc
SHLL [/etc] $ cd /
SHLL [/] $ pwd
/
SHLL [/] $ cd etc
SHLL [/etc] $ cd ..
SHLL [/] $ pwd
/

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_CD to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_CD when all shell commands have been configured.

PROGRAMMING INFORMATION:

The cd is implemented by a C language function which has the following prototype:

int rtems_shell_rtems_main_cd(
  int    argc,
  char **argv
);

The configuration structure for the cd has the following prototype:

extern rtems_shell_cmd_t rtems_shell_CD_Command;

4 Memory Commands

4.1 Introduction

The RTEMS shell has the following memory commands:

• mdump - Display contents of memory

• wdump - Display contents of memory (word)

• medit - Modify contents of memory

• mfill - File memory with pattern

• mmove - Move contents of memory

• malloc - Obtain information on C Program Heap

4.2 Commands

This section details the Memory Commands available. A subsection is dedicated to each of the commands and describes the behavior and configuration of that command as well as providing an example usage.

4.2.1 mdump - display contents of memory

SYNOPSYS:

     mdump [address [length]]

DESCRIPTION:

This command displays the contents of memory at the address and length in bytes specified on the command line.

When length is not provided, it defaults to 320 which is twenty lines of output with sixteen bytes of output per line.

When address is not provided, it defaults to 0x00000000.

EXIT STATUS:

This command always returns 0 to indicate success.

NOTES:

Dumping memory from a non-existent address may result in an unrecoverable program fault.

EXAMPLES:

The following is an example of how to use mdump:

      SHLL [/] $ mdump 0x10000 32
      0x0001000000 00 00 00 00 00 00 00-00 00 00 00 00 00 00 00 ................
      0x0001001000 00 00 00 00 00 00 00-00 00 00 00 00 00 00 00 ................
      SHLL [/] $ mdump 0x02000000 32
      0x02000000A1 48 00 00 29 00 80 33-81 C5 22 BC A6 10 21 00 .H..)..3.."...!.
      0x02000010A1 48 00 00 29 00 80 33-81 C5 22 BC A6 10 21 01 .H..)..3.."...!.
      SHLL [/] $ mdump 0x02001000 32
      0x0200100003 00 80 00 82 10 60 00-81 98 40 00 83 48 00 00 ......`.....H..
      0x0200101084 00 60 01 84 08 A0 07-86 10 20 01 87 28 C0 02 ..`....... ..(..

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_MDUMP to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_MDUMP when all shell commands have been configured.

PROGRAMMING INFORMATION:

The mdump is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_mdump(
       int    argc,
       char **argv
     );

The configuration structure for the mdump has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_MDUMP_Command;

4.2.2 wdump - display contents of memory (word)

SYNOPSYS:

     wdump [address [length]]

DESCRIPTION:

This command displays the contents of memory at the address and length in bytes specified on the command line.

When length is not provided, it defaults to 320 which is twenty lines of output with sixteen bytes of output per line.

When address is not provided, it defaults to 0x00000000.

EXIT STATUS:

This command always returns 0 to indicate success.

NOTES:

Dumping memory from a non-existent address may result in an unrecoverable program fault.

EXAMPLES:

The following is an example of how to use wdump:

      SHLL [/] $ wdump 0x02010000 32
      0x02010000 0201 08D8 0201 08C0-0201 08AC 0201 0874 ...............t
      0x02010010 0201 0894 0201 0718-0201 0640 0201 0798 ...............

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_WDUMP to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_WDUMP when all shell commands have been configured.

PROGRAMMING INFORMATION:

The wdump is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_wdump(
       int    argc,
       char **argv
     );

The configuration structure for the wdump has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_WDUMP_Command;

4.2.3 medit - modify contents of memory SYNOPSYS:

     medit address value1 [value2 ... valueN]

DESCRIPTION:

This command is used to modify the contents of the memory starting at address using the octets specified by the parameters value1 through valueN.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

Dumping memory from a non-existent address may result in an unrecoverable program fault.

EXAMPLES:

The following is an example of how to use medit:

      SHLL [/] $ mdump 0x02000000 32
      0x02000000 A1 48 00 00 29 00 80 33-81 C5 22 BC A6 10 21 00 .H..)..3.."...!.
      0x02000010 A1 48 00 00 29 00 80 33-81 C5 22 BC A6 10 21 01 .H..)..3.."...!.
      SHLL [/] $  medit 0x02000000 0x01 0x02 0x03 0x04 0x05 0x06 0x07 0x08 0x09
      SHLL [/] $ mdump 0x02000000 32
      0x02000000 01 02 03 04 05 06 07 08-09 00 22 BC A6 10 21 00 .........."...!.
      0x02000010 A1 48 00 00 29 00 80 33-81 C5 22 BC A6 10 21 01 .H..)..3.."...!.

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_MEDIT to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_MEDIT when all shell commands have been configured.

PROGRAMMING INFORMATION:

The medit is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_medit(
       int    argc,
       char **argv
     );

The configuration structure for the medit has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_MEDIT_Command;

4.2.4 mfill - file memory with pattern

SYNOPSYS:

     mfill address length value

DESCRIPTION:

This command is used to fill the memory starting at address for the specified length in octets when the specified at value.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

Filling a non-existent address range may result in an unrecoverable program fault. Similarly overwriting interrupt vector tables, code space or critical data areas can be fatal as shown in the example.

EXAMPLES:

In this example, the address used (0x23d89a0) as the base address of the filled area is the end of the stack for the Idle thread. This address was determined manually using gdb and is very specific to this application and BSP. The first command in this example is an mdump to display the initial contents of this memory. We see that the first 8 bytes are 0xA5 which is the pattern used as a guard by the Stack Checker. On the first context switch after the pattern is overwritten by the mfill command, the Stack Checker detect the pattern has been corrupted and generates a fatal error.

      SHLL [/] $ mdump 0x23d89a0 16
      0x023D89A0 A5 A5 A5 A5 A5 A5 A5 A5-FE ED F0 0D 0B AD 0D 06 ................
      SHLL [/] $ mfill 0x23d89a0 13 0x5a
      SHLL [/] $ BLOWN STACK!!! Offending task(0x23D4418): id=0x09010001; name=0x0203D908
        stack covers range 0x23D89A0 - 0x23D99AF (4112 bytes)
        Damaged pattern begins at 0x023D89A8 and is 16 bytes long

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_MFILL to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_MFILL when all shell commands have been configured.

PROGRAMMING INFORMATION:

The mfill is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_mfill(
       int    argc,
       char **argv
     );

The configuration structure for the mfill has the following prototype:

72 RTEMS Shell User’s Guide

     extern rtems_shell_cmd_t rtems_shell_MFILL_Command;

4.2.5 mmove - move contents of memory SYNOPSYS:

     mmove dst src length

DESCRIPTION:

This command is used to copy the contents of the memory starting at src to the memory located at dst for the specified length in octets.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use mmove:

      SHLL [/] $ mdump 0x023d99a0 16
      0x023D99A0 A5 A5 A5 A5 A5 A5 A5 A5-A5 A5 A5 A5 A5 A5 A5 A5 ................
      SHLL [/] $ mdump 0x02000000 16
      0x02000000 A1 48 00 00 29 00 80 33-81 C5 22 BC A6 10 21 00 .H..)..3.."...!.
      SHLL [/] $ mmove 0x023d99a0 0x02000000 13
      SHLL [/] $ mdump 0x023d99a0 16
      0x023D99A0 A1 48 00 00 29 00 80 33-81 C5 22 BC A6 A5 A5 A5 .H..)..3..".....

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_MMOVE to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_MMOVE when all shell commands have been configured.

PROGRAMMING INFORMATION:

The mmove is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_mmove(
       int    argc,
       char **argv
     );

The configuration structure for the mmove has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_MMOVE_Command;

4.2.6 malloc - obtain information on C program heap

SYNOPSYS:

 malloc [info|stats]

DESCRIPTION:

This command prints either information or statistics about the C Program Heap used by the malloc family of calls based upon the value of the first argument passed to the command. When the subcommand info is specified, information on the current state of the C Program Heap is reported. This includes the following information:

• Number of free blocks

• Largest free block

• Total bytes free

• Number of used blocks

• Largest used block

• Total bytes used

When the subcommand stats is specified, statistics on the the C Program Heap are reported. Malloc Family Statistics must be enabled for all of the values to be updated. The statistics available includes the following information:

•

• Currently available memory (in kilobytes)

• Currently allocated memory (in kilobytes)

• Maximum amount of memory ever allocated (in kilobytes)

• Lifetime tally of allocated memory (in kilobytes)

• Lifetime tally of freed memory (in kilobytes)

• Number of calls to malloc

• Number of calls to free

• Number of calls to realloc

• Number of calls to calloc

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

The CONFIGURE_MALLOC_STATISTICS confdefs.h constant must be defined when the ap-

plication is configured for the full set of statistics information to be available.

EXAMPLES:

The following is an example of how to use the malloc command.

SHLL [/] $ malloc info
Number of free blocks: 3
Largest free block:    3626672
Total bytes free:      3627768
Number of used blocks: 130
Largest used block:    1048
Total bytes used:      10136
SHLL [/] $ malloc stats
Malloc statistics
  avail:3552k  allocated:9k (0%) max:10k (0%) lifetime:21k freed:12k
  Call counts:   malloc:203   free:93   realloc:0   calloc:20
SHLL [/] $ malloc info
Number of free blocks: 3
Largest free block:    3626672
Total bytes free:      3627768
Number of used blocks: 130
Largest used block:    1048
Total bytes used:      10136
SHLL [/] $ malloc stats
Malloc statistics
  avail:3552k  allocated:9k (0%) max:10k (0%) lifetime:23k freed:14k
  Call counts:   malloc:205   free:95   realloc:0   calloc:20

Note that in the above example, the lifetime allocated and free values have increased between the two calls to malloc stats even though the amount of memory available in the C Program Heap is the same in both the malloc info invocations. This indicates that memory was allocated and freed as a side-effect of the commands.

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_MALLOC to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_MALLOC when all shell commands have been configured.

PROGRAMMING INFORMATION:

The malloc is implemented by a C language function which has the following prototype:

int rtems_shell_rtems_main_malloc(
  int    argc,
  char **argv
);

The configuration structure for the malloc has the following prototype:

extern rtems_shell_cmd_t rtems_shell_MALLOC_Command;

5 RTEMS Specific Commands

5.1 Introduction

The RTEMS shell has the following rtems commands:

• halt - Shutdown the system

• cpuuse - print or reset per thread cpu usage

• stackuse - print per thread stack usage

• perioduse - print or reset per period usage

• wkspace - Display information on Executive Workspace • config - Show the system configuration.

• itask - List init tasks for the system

• extension - Display information about extensions

• task - Display information about tasks

• queue - Display information about message queues

• sema - display information about semaphores

• region - display information about regions

• part - display information about partitions

• object - Display information about RTEMS objects • driver - Display the RTEMS device driver table

• dname - Displays information about named drivers

• pthread - Displays information about POSIX threads

5.2 Commands

This section details the RTEMS Specific Commands available. A subsection is dedicated to each of the commands and describes the behavior and configuration of that command as well as providing an example usage.

5.2.1 halt - Shutdown the system SYNOPSYS:

     halt

DESCRIPTION:

This command is used to shutdown the RTEMS application.

EXIT STATUS:

This command does not return.

NOTES: EXAMPLES:

The following is an example of how to use halt:

     SHLL [/] $ halt

The user will not see another prompt and the system will shutdown.

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_HALT to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_HALT when all shell commands have been configured.

PROGRAMMING INFORMATION:

The halt is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_halt(
       int    argc,
       char **argv
     );

The configuration structure for the halt has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_HALT_Command;

5.2.2 cpuuse - print or reset per thread cpu usage

SYNOPSYS:

     cpuuse [-r]

DESCRIPTION:

This command may be used to print a report on the per thread cpu usage or to reset the per thread CPU usage statistics. When invoked with the -r option, the CPU usage statistics are reset.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

The granularity of the timing information reported is dependent upon the BSP and the manner in which RTEMS was built. In the default RTEMS configuration, if the BSP supports nanosecond granularity timestamps, then the information reported will be highly accurate. Otherwise, the accuracy of the information reported is limited by the clock tick quantum.

EXAMPLES:

The following is an example of how to use cpuuse:

     SHLL [/] $ cpuuse
     CPU Usage by thread
        ID            NAME         SECONDS   PERCENT
     0x09010001   IDLE            49.745393   98.953
     0x0a010001   UI1              0.000000    0.000
     0x0a010002   SHLL             0.525928    1.046
     Time since last CPU Usage reset 50.271321 seconds
     SHLL [/] $ cpuuse -r
     Resetting CPU Usage information
     SHLL [/] $ cpuuse
     CPU Usage by thread
        ID            NAME         SECONDS   PERCENT
     0x09010001   IDLE             0.000000    0.000
     0x0a010001   UI1              0.000000    0.000
     0x0a010002   SHLL             0.003092  100.000
     Time since last CPU Usage reset 0.003092 seconds

In the above example, the system had set idle for nearly a minute when the first report was generated. The cpuuse -r and cpuuse commands were pasted from another window so were executed with no gap between. In the second report, only the shell thread has run since the CPU Usage was reset. It has consumed approximately 3.092 milliseconds of CPU time processing the two commands and generating the output.

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_CPUUSE to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_CPUUSE when all shell commands have been configured.

PROGRAMMING INFORMATION:

The cpuuse is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_cpuuse(
       int    argc,
       char **argv
     );

The configuration structure for the cpuuse has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_CPUUSE_Command;

5.2.3 stackuse - print per thread stack usage SYNOPSYS:

     stackuse

DESCRIPTION:

This command prints a Stack Usage Report for all of the tasks and threads in the system. On systems which support it, the usage of the interrupt stack is also included in the report.

EXIT STATUS:

This command always succeeds and returns 0.

NOTES:

The CONFIGURE_STACK_CHECKER_ENABLED confdefs.h constant must be defined when the

application is configured for this command to have any information to report.

EXAMPLES:

The following is an example of how to use stackuse:

      SHLL [/] $ stackuse
      Stack usage by thread
          ID      NAME    LOW          HIGH     CURRENT     AVAILABLE     USED
      0x09010001  IDLE 0x023d89a0 - 0x023d99af 0x023d9760      4096        608
      0x0a010001  UI1  0x023d9f30 - 0x023daf3f 0x023dad18      4096       1804
      0x0a010002  SHLL 0x023db4c0 - 0x023df4cf 0x023de9d0     16384       5116
      0xffffffff  INTR 0x023d2760 - 0x023d375f 0x00000000      4080        316

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_STACKUSE to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_STACKUSE when all shell commands have been configured.

PROGRAMMING INFORMATION:

The stackuse is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_stackuse(
       int    argc,
       char **argv
     );

The configuration structure for the stackuse has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_STACKUSE_Command;

5.2.4 perioduse - print or reset per period usage

SYNOPSYS:

     perioduse [-r]

DESCRIPTION:

This command may be used to print a statistics report on the rate monotonic periods in the application or to reset the rate monotonic period usage statistics. When invoked with the -r option, the usage statistics are reset.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

The granularity of the timing information reported is dependent upon the BSP and the manner in which RTEMS was built. In the default RTEMS configuration, if the BSP supports nanosecond granularity timestamps, then the information reported will be highly accurate. Otherwise, the accuracy of the information reported is limited by the clock tick quantum.

EXAMPLES:

The following is an example of how to use perioduse:

      SHLL [/] $ perioduse
      Period information by period
      --- CPU times are in seconds ---
      --- Wall times are in seconds ---
         ID     OWNER COUNT MISSED          CPU TIME                  WALL TIME
                                          MIN/MAX/AVG                MIN/MAX/AVG
      0x42010001 TA1    502      0 0:000039/0:042650/0:004158 0:000039/0:020118/0:002848
      0x42010002 TA2    502      0 0:000041/0:042657/0:004309 0:000041/0:020116/0:002848
      0x42010003 TA3    501      0 0:000041/0:041564/0:003653 0:000041/0:020003/0:002814
      0x42010004 TA4    501      0 0:000043/0:044075/0:004911 0:000043/0:020004/0:002814
      0x42010005 TA5     10      0 0:000065/0:005413/0:002739 0:000065/1:000457/0:041058
                                          MIN/MAX/AVG                MIN/MAX/AVG
      SHLL [/] $ perioduse -r
      Resetting Period Usage information
      SHLL [/] $ perioduse
      --- CPU times are in seconds ---
      --- Wall times are in seconds ---
         ID     OWNER COUNT MISSED          CPU TIME                  WALL TIME
                                          MIN/MAX/AVG                MIN/MAX/AVG
      0x42010001 TA1      0      0
      0x42010002 TA2      0      0
      0x42010003 TA3      0      0
      0x42010004 TA4      0      0
      0x42010005 TA5      0      0

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_PERIODUSE to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_PERIODUSE when all shell commands have been configured.

PROGRAMMING INFORMATION:

The perioduse is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_perioduse(
       int    argc,
       char **argv
     );

The configuration structure for the perioduse has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_PERIODUSE_Command;

5.2.5 wkspace - display information on executive workspace

SYNOPSYS:

     wkspace

DESCRIPTION:

This command prints information on the current state of the RTEMS Executive Workspace reported. This includes the following information:

• Number of free blocks

• Largest free block

• Total bytes free

• Number of used blocks

• Largest used block

• Total bytes used

EXIT STATUS:

This command always succeeds and returns 0.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use wkspace:

     SHLL [/] $ wkspace
     Number of free blocks: 1
     Largest free block:    132336
     Total bytes free:      132336
     Number of used blocks: 36
     Largest used block:    16408
     Total bytes used:      55344

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_WKSPACE to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_WKSPACE when all shell commands have been configured.

PROGRAMMING INFORMATION:

The wkspace is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_wkspace(
       int    argc,
       char **argv
     );

Chapter 5: RTEMS Specific Commands 85 The configuration structure for the wkspace has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_WKSPACE_Command;

5.2.6 config - show the system configuration. SYNOPSYS:

     config

DESCRIPTION:

This command display information about the RTEMS Configuration.

EXIT STATUS:

This command always succeeds and returns 0.

NOTES:

At this time, it does not report every configuration parameter. This is an area in which user submissions or sponsorship of a developer would be appreciated.

EXAMPLES:

The following is an example of how to use config:

      INITIAL (startup) Configuration Info
      ------------------------------------------------------------------------------
      WORKSPACE      start: 0x23d22e0;  size: 0x2dd20
      TIME           usec/tick: 10000;  tick/timeslice: 50;  tick/sec: 100
      MAXIMUMS       tasks: 20;  timers: 0;  sems: 50;  que's: 20;  ext's: 1
                     partitions: 0;  regions: 0;  ports: 0;  periods: 0

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_CONFIG to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_CONFIG when all shell commands have been configured.

PROGRAMMING INFORMATION:

The config is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_config(
       int    argc,
       char **argv
     );

The configuration structure for the config has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_CONFIG_Command;

5.2.7 itask - list init tasks for the system SYNOPSYS:

     itask

DESCRIPTION:

This command prints a report on the set of initialization tasks and threads in the system.

EXIT STATUS:

This command always succeeds and returns 0.

NOTES:

At this time, it includes only Classic API Initialization Tasks. This is an area in which user submissions or sponsorship of a developer would be appreciated.

EXAMPLES:

The following is an example of how to use itask:

      SHLL [/] $ itask
        #    NAME   ENTRY        ARGUMENT    PRIO   MODES  ATTRIBUTES   STACK SIZE
      ------------------------------------------------------------------------------
         0   UI1    [0x2002258] 0 [0x0]        1    nP      DEFAULT     4096 [0x1000]

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_ITASK to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_ITASK when all shell commands have been configured.

PROGRAMMING INFORMATION:

The itask is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_itask(
       int    argc,
       char **argv
     );

The configuration structure for the itask has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_ITASK_Command;

5.2.8 extension - display information about extensions

SYNOPSYS:

     extension [id [id ...] ]

DESCRIPTION:

When invoked with no arguments, this command prints information on the set of User Extensions currently active in the system.

If invoked with a set of ids as arguments, then just those objects are included in the information printed.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of using the extension command on a system with no user extensions.

      SHLL [/] $ extension
        ID       NAME
      ------------------------------------------------------------------------------

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_EXTENSION to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_EXTENSION when all shell commands have been configured.

PROGRAMMING INFORMATION:

The extension is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_extension(
       int    argc,
       char **argv
     );

The configuration structure for the extension has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_EXTENSION_Command;

5.2.9 task - display information about tasks

SYNOPSYS:

     task [id [id ...] ]

DESCRIPTION:

When invoked with no arguments, this command prints information on the set of Classic API Tasks currently active in the system.

If invoked with a set of ids as arguments, then just those objects are included in the information printed.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use the task on an application with just two Classic API tasks:

      SHLL [/] $ task
        ID       NAME   PRIO   STAT   MODES  EVENTS   WAITID  WAITARG  NOTES
      ------------------------------------------------------------------------------
      0a010001   UI1      1   SUSP   P:T:nA  NONE
      0a010002   SHLL   100   READY  P:T:nA  NONE

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_TASK to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_TASK when all shell commands have been configured.

PROGRAMMING INFORMATION:

The task is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_task(
       int    argc,
       char **argv
     );

The configuration structure for the task has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_TASK_Command;

5.2.10 queue - display information about message queues

SYNOPSYS:

     queue [id [id ... ] ]

DESCRIPTION:

When invoked with no arguments, this command prints information on the set of Classic API Message Queues currently active in the system.

If invoked with a set of ids as arguments, then just those objects are included in the information printed.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of using the queue command on a system with no Classic API Message Queues.

      SHLL [/] $ queue
        ID       NAME   ATTRIBUTES   PEND   MAXPEND  MAXSIZE
      ------------------------------------------------------------------------------

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_QUEUE to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_QUEUE when all shell commands have been configured.

PROGRAMMING INFORMATION:

The queue is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_queue(
       int    argc,
       char **argv
     );

The configuration structure for the queue has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_QUEUE_Command;

5.2.11 sema - display information about semaphores

SYNOPSYS:

     sema [id [id ... ] ]

DESCRIPTION:

When invoked with no arguments, this command prints information on the set of Classic API Semaphores currently active in the system.

If invoked with a set of objects ids as arguments, then just those objects are included in the information printed.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use sema:

      SHLL [/] $ sema
        ID       NAME   ATTR        PRICEIL CURR_CNT HOLDID
      ------------------------------------------------------------------------------
      1a010001   LBIO   PR:BI:IN      0        1     00000000
      1a010002   TRmi   PR:BI:IN      0        1     00000000
      1a010003   LBI00  PR:BI:IN      0        1     00000000
      1a010004   TRia   PR:BI:IN      0        1     00000000
      1a010005   TRoa   PR:BI:IN      0        1     00000000
      1a010006   TRxa   <assoc.c: BAD NAME>   0    0 09010001
      1a010007   LBI01  PR:BI:IN      0        1     00000000
      1a010008   LBI02  PR:BI:IN      0        1     00000000

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_SEMA to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_SEMA when all shell commands have been configured.

PROGRAMMING INFORMATION:

The sema is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_sema(
       int    argc,
       char **argv
     );

The configuration structure for the sema has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_SEMA_Command;

5.2.12 region - display information about regions

SYNOPSYS:

     region [id [id ... ] ]

DESCRIPTION:

When invoked with no arguments, this command prints information on the set of Classic API Regions currently active in the system.

If invoked with a set of object ids as arguments, then just those object are included in the information printed.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of using the region command on a system with no user extensions.

      SHLL [/] $ region
        ID       NAME   ATTR        STARTADDR LENGTH    PAGE_SIZE USED_BLOCKS
      ------------------------------------------------------------------------------

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_REGION to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_REGION when all shell commands have been configured.

PROGRAMMING INFORMATION:

The region is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_region(
       int    argc,
       char **argv
     );

The configuration structure for the region has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_REGION_Command;

5.2.13 part - display information about partitions SYNOPSYS:

     part [id [id ... ] ]

DESCRIPTION:

When invoked with no arguments, this command prints information on the set of Classic API Partitions currently active in the system. If invoked with a set of object ids as arguments, then just those objects are included in the information printed.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of using the part command on a system with no user extensions.

      SHLL [/] $ part
        ID       NAME   ATTR        STARTADDR LENGTH    BUF_SIZE  USED_BLOCKS
      ------------------------------------------------------------------------------

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_PART to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_PART when all shell commands have been configured.

PROGRAMMING INFORMATION:

The part is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_part(
       int    argc,
       char **argv
     );

The configuration structure for the part has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_PART_Command;

5.2.14 object - display information about rtems objects SYNOPSYS:

     object [id [id ...] ]

DESCRIPTION:

When invoked with a set of object ids as arguments, then a report on those objects is printed.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use object:

      SHLL [/] $ object 0a010001 1a010002
        ID       NAME   PRIO   STAT   MODES  EVENTS   WAITID  WAITARG  NOTES
      ------------------------------------------------------------------------------
      0a010001   UI1      1   SUSP   P:T:nA  NONE
        ID       NAME   ATTR        PRICEIL CURR_CNT HOLDID
      ------------------------------------------------------------------------------
      1a010002   TRmi   PR:BI:IN      0        1     00000000

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_OBJECT to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_OBJECT when all shell commands have been configured.

PROGRAMMING INFORMATION:

The object is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_object(
       int    argc,
       char **argv
     );

The configuration structure for the object has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_OBJECT_Command;

5.2.15 driver - display the rtems device driver table

SYNOPSYS:

     driver [ major [ major ... ] ]

DESCRIPTION:

When invoked with no arguments, this command prints information on the set of Device Drivers currently active in the system.

If invoked with a set of major numbers as arguments, then just those Device Drivers are included in the information printed.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use driver:

      SHLL [/] $ driver
        Major      Entry points
      ------------------------------------------------------------------------------
        0          init: [0x200256c];  control: [0x20024c8]
                   open: [0x2002518];  close: [0x2002504]
                   read: [0x20024f0];  write: [0x20024dc]
        1          init: [0x20023fc];  control: [0x2002448]
                   open: [0x0];  close: [0x0]
                   read: [0x0];  write: [0x0]
      SHLL [/] $

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_DRIVER to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_DRIVER when all shell commands have been configured.

PROGRAMMING INFORMATION:

The driver is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_driver(
       int    argc,
       char **argv
     );

The configuration structure for the driver has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_DRIVER_Command;

5.2.16 dname - displays information about named drivers SYNOPSYS:

     dname

DESCRIPTION:

This command XXX WARNING! XXX This command does not appear to work as of 27 February 2008.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use dname:

     EXAMPLE_TBD

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_DNAME to have this command included. This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_DNAME when all shell commands have been configured.

PROGRAMMING INFORMATION:

The dname is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_dname(
       int    argc,
       char **argv
     );

The configuration structure for the dname has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_DNAME_Command;

5.2.17 pthread - display information about POSIX threads SYNOPSYS:

     pthread [id [id ...] ]

DESCRIPTION:

When invoked with no arguments, this command prints information on the set of POSIX API threads currently active in the system. If invoked with a set of ids as arguments, then just those objects are included in the information printed.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

This command is only available when the POSIX API is configured.

EXAMPLES:

The following is an example of how to use the task on an application with four POSIX threads:

      SHLL [/] $ pthread
        ID       NAME           PRI  STATE MODES   EVENTS    WAITID  WAITARG  NOTES
      ------------------------------------------------------------------------------
      0b010002   Main           133 READY  P:T:nA    NONE   43010001 0x7b1148
      0b010003   ISR            133 Wcvar  P:T:nA    NONE   43010003 0x7b1148
      0b01000c                  133 READY  P:T:nA    NONE   33010002 0x7b1148
      0b01000d                  133 Wmutex P:T:nA    NONE   33010002 0x7b1148

CONFIGURATION:

This command is part of the monitor commands which are always available in the shell.

PROGRAMMING INFORMATION:

This command is not directly available for invocation.

6 Network Commands

6.1 Introduction

The RTEMS shell has the following network commands:

• netstats - obtain network statistics

• ifconfig - configure a network interface

• route - show or manipulate the IP routing table

6.2 Commands

This section details the Network Commands available. A subsection is dedicated to each of the commands and describes the behavior and configuration of that command as well as providing an example usage.

6.2.1 netstats - obtain network statistics

SYNOPSYS:

     netstats [-Aimfpcut]

DESCRIPTION:

This command is used to display various types of network statistics. The information displayed can be specified using command line arguments in various combinations. The arguments are interpreted as follows:

-A print All statistics

-i print Inet Routes

-m print MBUF Statistics

-f print IF Statistics

-p print IP Statistics

-c print ICMP Statistics

-u print UDP Statistics

-t print TCP Statistics

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

NONE

EXAMPLES:

The following is an example of how to use netstats:

The following is an example of using the netstats command to print the IP routing table:

      [/] $ netstats -i
      Destination     Gateway/Mask/Hw    Flags     Refs     Use Expire Interface
      default         192.168.1.14       UGS         0        0      0 eth1
      192.168.1.0     255.255.255.0      U           0        0      1 eth1
      192.168.1.14    00:A0:C8:1C:EE:28  UHL         1        0   1219 eth1
      192.168.1.51    00:1D:7E:0C:D0:7C  UHL         0      840   1202 eth1
      192.168.1.151   00:1C:23:B2:0F:BB  UHL         1       23   1219 eth1

The following is an example of using the netstats command to print the MBUF statistics:

      [/] $ netstats -m
      ************ MBUF STATISTICS ************
      mbufs:2048    clusters: 128    free:  63
      drops:   0       waits:   0  drains:   0
            free:1967          data:79          header:2           socket:0
             pcb:0           rtable:0           htable:0           atable:0
          soname:0           soopts:0           ftable:0           rights:0
          ifaddr:0          control:0          oobdata:0

Chapter 6: Network Commands 101

The following is an example of using the netstats command to print the print the interface statistics:

[/] $ netstats -f
************ INTERFACE STATISTICS ************
***** eth1 *****
Ethernet Address: 00:04:9F:00:5B:21
Address:192.168.1.244   Broadcast Address:192.168.1.255   Net mask:255.255.255.0
Flags: Up Broadcast Running Active Multicast
Send queue limit:50   length:1    Dropped:0
      Rx Interrupts:889            Not First:0               Not Last:0
              Giant:0              Non-octet:0
            Bad CRC:0                Overrun:0              Collision:0
      Tx Interrupts:867             Deferred:0         Late Collision:0
   Retransmit Limit:0               Underrun:0             Misaligned:0

The following is an example of using the netstats command to print the print IP statistics:

[/] $ netstats -p
************ IP Statistics ************
             total packets received         894
  packets rcvd for unreachable dest          13
 datagrams delivered to upper level         881
    total ip packets generated here         871

The following is an example of using the netstats command to print the ICMP statistics:

[/] $ netstats -c
************ ICMP Statistics ************
                        Type 0 sent         843
                number of responses         843
                    Type 8 received         843

The following is an example of using the netstats command to print the UDP statistics:

[/] $ netstats -u
************ UDP Statistics ************

The following is an example of using the netstats command to print the TCP statistics:

[/] $ netstats -t
************ TCP Statistics ************
               connections accepted           1
            connections established           1
     segs where we tried to get rtt          34
                 times we succeeded          35
                  delayed acks sent           2
                 total packets sent          37
                  data packets sent          35
                    data bytes sent        2618
              ack-only packets sent           2
             total packets received          47
       packets received in sequence          12
         bytes received in sequence         307
                   rcvd ack packets          35
           bytes acked by rcvd acks        2590
      times hdr predict ok for acks          27
 times hdr predict ok for data pkts          10

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_NETSTATS to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_NETSTATS when all shell commands have been configured.

PROGRAMMING INFORMATION:

The netstats is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_netstats(
       int    argc,
       char **argv
     );

The configuration structure for the netstats has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_NETSTATS_Command;

6.2.2 ifconfig - configure a network interface

SYNOPSYS:

     ifconfig
     ifconfig interface
     ifconfig interface [up|down]
     ifconfig interface [netmask|pointtopoint|broadcast] IP

DESCRIPTION:

This command may be used to display information about the network interfaces in the system or configure them.

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

Just like its counterpart on GNU/Linux and BSD systems, this command is complicated. More example usages would be a welcome submission.

EXAMPLES:

The following is an example of how to use ifconfig:

       ************ INTERFACE STATISTICS ************
      ***** eth1 *****
      Ethernet Address: 00:04:9F:00:5B:21
      Address:192.168.1.244   Broadcast Address:192.168.1.255   Net mask:255.255.255.0
      Flags: Up Broadcast Running Active Multicast
      Send queue limit:50   length:1    Dropped:0
            Rx Interrupts:5391           Not First:0               Not Last:0
                    Giant:0              Non-octet:0
                  Bad CRC:0                Overrun:0              Collision:0
            Tx Interrupts:5256            Deferred:0         Late Collision:0
         Retransmit Limit:0               Underrun:0             Misaligned:0

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_IFCONFIG to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_IFCONFIG when all shell commands have been configured.

PROGRAMMING INFORMATION:

The ifconfig is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_ifconfig(
       int    argc,
       char **argv
     );

The configuration structure for the ifconfig has the following prototype:

104 RTEMS Shell User’s Guide

     extern rtems_shell_cmd_t rtems_shell_IFCONFIG_Command;

Chapter 6: Network Commands 105

6.2.3 route - show or manipulate the ip routing table

SYNOPSYS:

route [subcommand] [args]

DESCRIPTION:

This command is used to display and manipulate the routing table. When invoked with no arguments, the current routing information is displayed. When invoked with the subcommands add or del, then additional arguments must be provided to describe the route.

Command templates include the following:

route [add|del] -net IP_ADDRESS gw GATEWAY_ADDRESS [netmask MASK]
route [add|del] -host IP_ADDRESS gw GATEWAY_ADDRES [netmask MASK]

When not provided the netmask defaults to 255.255.255.0

EXIT STATUS:

This command returns 0 on success and non-zero if an error is encountered.

NOTES:

Just like its counterpart on GNU/Linux and BSD systems, this command is complicated. More example usages would be a welcome submission.

EXAMPLES:

The following is an example of how to use route to display, add, and delete a new route:

[/] $ route
Destination     Gateway/Mask/Hw    Flags     Refs     Use Expire Interface
default         192.168.1.14       UGS         0        0      0 eth1
192.168.1.0     255.255.255.0      U           0        0      1 eth1
192.168.1.14    00:A0:C8:1C:EE:28  UHL         1        0   1444 eth1
192.168.1.51    00:1D:7E:0C:D0:7C  UHL         0    10844   1202 eth1
192.168.1.151   00:1C:23:B2:0F:BB  UHL         2       37   1399 eth1
[/] $  route add -net 192.168.3.0 gw 192.168.1.14
[/] $ route
Destination     Gateway/Mask/Hw    Flags     Refs     Use Expire Interface
default         192.168.1.14       UGS         0        0      0 eth1
192.168.1.0     255.255.255.0      U           0        0      1 eth1
192.168.1.14    00:A0:C8:1C:EE:28  UHL         2        0   1498 eth1
192.168.1.51    00:1D:7E:0C:D0:7C  UHL         0    14937   1202 eth1
192.168.1.151   00:1C:23:B2:0F:BB  UHL         2       96   1399 eth1
192.168.3.0     192.168.1.14       UGS         0        0      0 eth1
[/] $ route del -net 192.168.3.0 gw 192.168.1.14
[/] $ route
Destination     Gateway/Mask/Hw    Flags     Refs     Use Expire Interface
default         192.168.1.14       UGS         0        0      0 eth1
192.168.1.0     255.255.255.0      U           0        0      1 eth1
192.168.1.14    00:A0:C8:1C:EE:28  UHL         1        0   1498 eth1
192.168.1.51    00:1D:7E:0C:D0:7C  UHL         0    15945   1202 eth1
192.168.1.151   00:1C:23:B2:0F:BB  UHL         2      117   1399 eth1

CONFIGURATION:

This command is included in the default shell command set. When building a custom command set, define CONFIGURE_SHELL_COMMAND_ROUTE to have this command included.

This command can be excluded from the shell command set by defining CONFIGURE_SHELL_ NO_COMMAND_ROUTE when all shell commands have been configured.

PROGRAMMING INFORMATION:

The route is implemented by a C language function which has the following prototype:

     int rtems_shell_rtems_main_route(
       int    argc,
       char **argv
     );

The configuration structure for the route has the following prototype:

     extern rtems_shell_cmd_t rtems_shell_ROUTE_Command;

Function and Variable Index 107

Function and Variable Index

C

CONFIGURE_MALLOC_STATISTICS . . . . . . . . . . . . . . . . 74
CONFIGURE_SHELL_COMMAND_ALIAS . . . . . . . . . . . . . . . 8
CONFIGURE_SHELL_COMMAND_BLKSYNC . . . . . . . . . . . 49
CONFIGURE_SHELL_COMMAND_CAT . . . . . . . . . . . . . . . . 44
CONFIGURE_SHELL_COMMAND_CD . . . . . . . . . . . . . . . . . 65
CONFIGURE_SHELL_COMMAND_CHDIR . . . . . . . . . . . . . . 33
CONFIGURE_SHELL_COMMAND_CHMOD . . . . . . . . . . . . . . 43
CONFIGURE_SHELL_COMMAND_CHROOT. . . . . . . . . . . . . 41
CONFIGURE_SHELL_COMMAND_CONFIG. . . . . . . . . . . . . 86
CONFIGURE_SHELL_COMMAND_CP . . . . . . . . . . . . . . . . . 27
CONFIGURE_SHELL_COMMAND_CPUUSE. . . . . . . . . . . . . 80
CONFIGURE_SHELL_COMMAND_DATE . . . . . . . . . . . . . . . . 9
CONFIGURE_SHELL_COMMAND_DD . . . . . . . . . . . . . . . . . 53
CONFIGURE_SHELL_COMMAND_DEBUGRFS . . . . . . . . . . 63
CONFIGURE_SHELL_COMMAND_DIR . . . . . . . . . . . . . . . . 60
CONFIGURE_SHELL_COMMAND_DNAME . . . . . . . . . . . . . . 96
CONFIGURE_SHELL_COMMAND_DRIVER. . . . . . . . . . . . . 95
CONFIGURE_SHELL_COMMAND_ECHO . . . . . . . . . . . . . . . 11
CONFIGURE_SHELL_COMMAND_EXTENSION . . . . . . . . . 88
CONFIGURE_SHELL_COMMAND_FDISK . . . . . . . . . . . . . . 59
CONFIGURE_SHELL_COMMAND_GETENV. . . . . . . . . . . . . 16
CONFIGURE_SHELL_COMMAND_HALT . . . . . . . . . . . . . . . 78
CONFIGURE_SHELL_COMMAND_HEXDUMP . . . . . . . . . . . 58
CONFIGURE_SHELL_COMMAND_ID . . . . . . . . . . . . . . . . . 13
CONFIGURE_SHELL_COMMAND_IFCONFIG . . . . . . . . . 103
CONFIGURE_SHELL_COMMAND_ITASK . . . . . . . . . . . . . . 87
CONFIGURE_SHELL_COMMAND_LN . . . . . . . . . . . . . . . . . 38
CONFIGURE_SHELL_COMMAND_LOGOFF. . . . . . . . . . . . . 20
CONFIGURE_SHELL_COMMAND_LS . . . . . . . . . . . . . . . . . 31
CONFIGURE_SHELL_COMMAND_MALLOC. . . . . . . . . . . . . 75
CONFIGURE_SHELL_COMMAND_MDUMP . . . . . . . . . . . . . . 68
CONFIGURE_SHELL_COMMAND_MEDIT . . . . . . . . . . . . . . 70
CONFIGURE_SHELL_COMMAND_MFILL . . . . . . . . . . . . . . 71
CONFIGURE_SHELL_COMMAND_MKDIR . . . . . . . . . . . . . . 34
CONFIGURE_SHELL_COMMAND_MKNOD . . . . . . . . . . . . . . 40
CONFIGURE_SHELL_COMMAND_MKRFS . . . . . . . . . . . . . . 61
CONFIGURE_SHELL_COMMAND_MMOVE . . . . . . . . . . . . . . 73
CONFIGURE_SHELL_COMMAND_MOUNT . . . . . . . . . . . . . . 46
CONFIGURE_SHELL_COMMAND_MV . . . . . . . . . . . . . . . . . 29
CONFIGURE_SHELL_COMMAND_NETSTATS . . . . . . . . . 102
CONFIGURE_SHELL_COMMAND_OBJECT. . . . . . . . . . . . . 94
CONFIGURE_SHELL_COMMAND_PART . . . . . . . . . . . . . . . 93
CONFIGURE_SHELL_COMMAND_PERIODUSE . . . . . . . . . 82
CONFIGURE_SHELL_COMMAND_PWD . . . . . . . . . . . . . . . . 30
CONFIGURE_SHELL_COMMAND_QUEUE . . . . . . . . . . . . . . 90
CONFIGURE_SHELL_COMMAND_REGION. . . . . . . . . . . . . 92
CONFIGURE_SHELL_COMMAND_RM . . . . . . . . . . . . . . . . . 45
CONFIGURE_SHELL_COMMAND_RMDIR . . . . . . . . . . . . . . 36
CONFIGURE_SHELL_COMMAND_ROUTE. . . . . . . . . . . . . 106
CONFIGURE_SHELL_COMMAND_RTC . . . . . . . . . . . . . . . . 21
CONFIGURE_SHELL_COMMAND_SEMA . . . . . . . . . . . . . . . 91
CONFIGURE_SHELL_COMMAND_SETENV. . . . . . . . . . . . . 17
CONFIGURE_SHELL_COMMAND_SLEEP . . . . . . . . . . . . . . 12

CONFIGURE_SHELL_COMMAND_STACKUSE . . . . . . . . . . 81
CONFIGURE_SHELL_COMMAND_TASK . . . . . . . . . . . . . . . 89
CONFIGURE_SHELL_COMMAND_TIME . . . . . . . . . . . . . . . 19
CONFIGURE_SHELL_COMMAND_TTY . . . . . . . . . . . . . . . . 14
CONFIGURE_SHELL_COMMAND_UMASK . . . . . . . . . . . . . . 24
CONFIGURE_SHELL_COMMAND_UNMOUNT . . . . . . . . . . . 48
CONFIGURE_SHELL_COMMAND_UNSETENV . . . . . . . . . . 18
CONFIGURE_SHELL_COMMAND_WDUMP . . . . . . . . . . . . . . 69
CONFIGURE_SHELL_COMMAND_WHOAMI. . . . . . . . . . . . . 15
CONFIGURE_SHELL_COMMAND_WKSPACE . . . . . . . . . . . 84
CONFIGURE_SHELL_NO_COMMAND_ALIAS . . . . . . . . . . . 8
CONFIGURE_SHELL_NO_COMMAND_BLKSYNC . . . . . . . . 49
CONFIGURE_SHELL_NO_COMMAND_CAT. . . . . . . . . . . . . 44
CONFIGURE_SHELL_NO_COMMAND_CD . . . . . . . . . . . . . . 65
CONFIGURE_SHELL_NO_COMMAND_CHDIR . . . . . . . . . . 33
CONFIGURE_SHELL_NO_COMMAND_CHMOD . . . . . . . . . . 43
CONFIGURE_SHELL_NO_COMMAND_CHROOT . . . . . . . . . 41
CONFIGURE_SHELL_NO_COMMAND_CONFIG . . . . . . . . . 86
CONFIGURE_SHELL_NO_COMMAND_CP . . . . . . . . . . . . . . 27
CONFIGURE_SHELL_NO_COMMAND_CPUUSE . . . . . . . . . 80
CONFIGURE_SHELL_NO_COMMAND_DATE. . . . . . . . . . . . . 9
CONFIGURE_SHELL_NO_COMMAND_DD . . . . . . . . . . . . . . 53
CONFIGURE_SHELL_NO_COMMAND_DEBUGRFS . . . . . . . 63
CONFIGURE_SHELL_NO_COMMAND_DIR. . . . . . . . . . . . . 60
CONFIGURE_SHELL_NO_COMMAND_DNAME . . . . . . . . . . 96
CONFIGURE_SHELL_NO_COMMAND_DRIVER . . . . . . . . . 95
CONFIGURE_SHELL_NO_COMMAND_ECHO . . . . . . . . . . . 11
CONFIGURE_SHELL_NO_COMMAND_EXTENSION. . . . . . 88
CONFIGURE_SHELL_NO_COMMAND_FDISK . . . . . . . . . . 59
CONFIGURE_SHELL_NO_COMMAND_GETENV . . . . . . . . . 16
CONFIGURE_SHELL_NO_COMMAND_HALT . . . . . . . . . . . 78
CONFIGURE_SHELL_NO_COMMAND_HEXDUMP . . . . . . . . 58
CONFIGURE_SHELL_NO_COMMAND_ID . . . . . . . . . . . . . . 13
CONFIGURE_SHELL_NO_COMMAND_IFCONFIG . . . . . . 103
CONFIGURE_SHELL_NO_COMMAND_ITASK . . . . . . . . . . 87
CONFIGURE_SHELL_NO_COMMAND_LN . . . . . . . . . . . . . . 38
CONFIGURE_SHELL_NO_COMMAND_LOGOFF . . . . . . . . . 20
CONFIGURE_SHELL_NO_COMMAND_LS . . . . . . . . . . . . . . 31
CONFIGURE_SHELL_NO_COMMAND_MALLOC . . . . . . . . . 75
CONFIGURE_SHELL_NO_COMMAND_MDUMP . . . . . . . . . . 68
CONFIGURE_SHELL_NO_COMMAND_MEDIT . . . . . . . . . . 70
CONFIGURE_SHELL_NO_COMMAND_MFILL . . . . . . . . . . 71
CONFIGURE_SHELL_NO_COMMAND_MKDIR . . . . . . . . . . 34
CONFIGURE_SHELL_NO_COMMAND_MKNOD . . . . . . . . . . 40
CONFIGURE_SHELL_NO_COMMAND_MKRFS . . . . . . . . . . 61
CONFIGURE_SHELL_NO_COMMAND_MMOVE . . . . . . . . . . 73
CONFIGURE_SHELL_NO_COMMAND_MOUNT . . . . . . . . . . 46
CONFIGURE_SHELL_NO_COMMAND_MV . . . . . . . . . . . . . . 29
CONFIGURE_SHELL_NO_COMMAND_NETSTATS . . . . . . 102
CONFIGURE_SHELL_NO_COMMAND_OBJECT . . . . . . . . . 94
CONFIGURE_SHELL_NO_COMMAND_PART . . . . . . . . . . . 93
CONFIGURE_SHELL_NO_COMMAND_PERIODUSE. . . . . . 82
CONFIGURE_SHELL_NO_COMMAND_PWD. . . . . . . . . . . . . 30
CONFIGURE_SHELL_NO_COMMAND_QUEUE . . . . . . . . . . 90
CONFIGURE_SHELL_NO_COMMAND_REGION . . . . . . . . . 92

108 RTEMS Shell User’s Guide

CONFIGURE_SHELL_NO_COMMAND_RM . . . . . . . . . . . . . . 45
CONFIGURE_SHELL_NO_COMMAND_RMDIR . . . . . . . . . . 36
CONFIGURE_SHELL_NO_COMMAND_ROUTE . . . . . . . . . 106
CONFIGURE_SHELL_NO_COMMAND_RTC. . . . . . . . . . . . . 21
CONFIGURE_SHELL_NO_COMMAND_SEMA . . . . . . . . . . . 91
CONFIGURE_SHELL_NO_COMMAND_SETENV . . . . . . . . . 17
CONFIGURE_SHELL_NO_COMMAND_SLEEP . . . . . . . . . . 12
CONFIGURE_SHELL_NO_COMMAND_STACKUSE . . . . . . . 81
CONFIGURE_SHELL_NO_COMMAND_TASK . . . . . . . . . . . 89
CONFIGURE_SHELL_NO_COMMAND_TIME . . . . . . . . . . . 19
CONFIGURE_SHELL_NO_COMMAND_TTY. . . . . . . . . . . . . 14
CONFIGURE_SHELL_NO_COMMAND_UMASK . . . . . . . . . . 24
CONFIGURE_SHELL_NO_COMMAND_UNMOUNT . . . . . . . . 48
CONFIGURE_SHELL_NO_COMMAND_UNSETENV . . . . . . . 18
CONFIGURE_SHELL_NO_COMMAND_WDUMP . . . . . . . . . . 69
CONFIGURE_SHELL_NO_COMMAND_WHOAMI . . . . . . . . . 15
CONFIGURE_SHELL_NO_COMMAND_WKSPACE . . . . . . . . 84

R

rtems_shell_init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

rtems_shell_rtems_main_alias . . . . . . . . . . . . . . . . 8
rtems_shell_rtems_main_blksync. . . . . . . . . . . . . 49
rtems_shell_rtems_main_cat . . . . . . . . . . . . . . . . . 44
rtems_shell_rtems_main_cd . . . . . . . . . . . . . . . . . . 65
rtems_shell_rtems_main_chdir . . . . . . . . . . . . . . . 33
rtems_shell_rtems_main_chmod . . . . . . . . . . . . . . . 43
rtems_shell_rtems_main_chroot . . . . . . . . . . . . . . 41
rtems_shell_rtems_main_config . . . . . . . . . . . . . . 86
rtems_shell_rtems_main_cp . . . . . . . . . . . . . . . . . . 27
rtems_shell_rtems_main_cpuuse . . . . . . . . . . . . . . 80
rtems_shell_rtems_main_date . . . . . . . . . . . . . . . . . 9
rtems_shell_rtems_main_dd . . . . . . . . . . . . . . . . . . 53
rtems_shell_rtems_main_debugrfs . . . . . . . . . . . 64
rtems_shell_rtems_main_dir . . . . . . . . . . . . . . . . . 60
rtems_shell_rtems_main_dname . . . . . . . . . . . . . . . 96
rtems_shell_rtems_main_driver . . . . . . . . . . . . . . 95
rtems_shell_rtems_main_echo . . . . . . . . . . . . . . . . 11
rtems_shell_rtems_main_extension . . . . . . . . . . 88
rtems_shell_rtems_main_getenv . . . . . . . . . . . . . . 16

rtems_shell_rtems_main_halt . . . . . . . . . . . . . . . . 78
rtems_shell_rtems_main_hexdump. . . . . . . . . . . . . 58
rtems_shell_rtems_main_id . . . . . . . . . . . . . . . . . . 13
rtems_shell_rtems_main_ifconfig . . . . . . . . . . 103
rtems_shell_rtems_main_itask . . . . . . . . . . . . . . . 87
rtems_shell_rtems_main_ln . . . . . . . . . . . . . . . . . . 38
rtems_shell_rtems_main_logoff . . . . . . . . . . . . . . 20
rtems_shell_rtems_main_ls . . . . . . . . . . . . . . . . . . 31
rtems_shell_rtems_main_malloc . . . . . . . . . . . . . . 75
rtems_shell_rtems_main_mdump . . . . . . . . . . . . . . . 68
rtems_shell_rtems_main_medit . . . . . . . . . . . . . . . 70
rtems_shell_rtems_main_mfill . . . . . . . . . . . . . . . 71
rtems_shell_rtems_main_mkdir . . . . . . . . . . . . . . . 34
rtems_shell_rtems_main_mknod . . . . . . . . . . . . . . . 40
rtems_shell_rtems_main_mkrfs . . . . . . . . . . . . . . . 62
rtems_shell_rtems_main_mmove . . . . . . . . . . . . . . . 73
rtems_shell_rtems_main_mount . . . . . . . . . . . . . . . 47
rtems_shell_rtems_main_mv . . . . . . . . . . . . . . . . . . 29
rtems_shell_rtems_main_netstats . . . . . . . . . . 102
rtems_shell_rtems_main_object . . . . . . . . . . . . . . 94
rtems_shell_rtems_main_part . . . . . . . . . . . . . . . . 93
rtems_shell_rtems_main_perioduse . . . . . . . . . . 83
rtems_shell_rtems_main_pwd . . . . . . . . . . . . . . . . . 30
rtems_shell_rtems_main_queue . . . . . . . . . . . . . . . 90
rtems_shell_rtems_main_region . . . . . . . . . . . . . . 92
rtems_shell_rtems_main_rm . . . . . . . . . . . . . . . . . . 45
rtems_shell_rtems_main_rmdir . . . . . . . . . . . . . . . 36
rtems_shell_rtems_main_route . . . . . . . . . . . . . . 106
rtems_shell_rtems_main_sema . . . . . . . . . . . . . . . . 91
rtems_shell_rtems_main_setenv . . . . . . . . . . . . . . 17
rtems_shell_rtems_main_sleep . . . . . . . . . . . . . . . 12
rtems_shell_rtems_main_stackuse . . . . . . . . . . . 81
rtems_shell_rtems_main_task . . . . . . . . . . . . . . . . 89
rtems_shell_rtems_main_time . . . . . . . . . . . . . . . . 19
rtems_shell_rtems_main_tty . . . . . . . . . . . . . . . . . 14
rtems_shell_rtems_main_umask . . . . . . . . . . . . . . . 24
rtems_shell_rtems_main_unmount. . . . . . . . . . . . . 48
rtems_shell_rtems_main_unsetenv . . . . . . . . . . . 18
rtems_shell_rtems_main_wdump . . . . . . . . . . . . . . . 69
rtems_shell_rtems_main_whoami . . . . . . . . . . . . . . 15
rtems_shell_rtems_main_wkspace. . . . . . . . . . . . . 84

Concept Index 109

Concept Index

initialization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Command Index 111

Command Index

A

alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

B

blksync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

C

cat. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 cd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 chdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 chmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 chroot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 config . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 cp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 cpuuse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

D

date. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 dd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 debugrfs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 dname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 driver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

E

echo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 extension . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88

F

fdisk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

G

getenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

H

halt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 hexdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

I

id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ifconfig . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 itask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

L

ln . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 logoff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

M

malloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 mdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 medit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 mfill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 mkdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 mknod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 mkrfs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 mmove . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 mount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 mv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

N

netstats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

O

object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

P

part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 perioduse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 pthread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 pwd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

Q

queue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

R

region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 rm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 rmdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 route . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 rtc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

S

sema . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 setenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 stackuse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

112 RTEMS Shell User’s Guide

T

task . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 tty. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

U

umask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 unmount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 unsetenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

W

wdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 whoami . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 wkspace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84